### v2

# Mora: Enabling Generalist Video Generation via A Multi-Agent Framework

Abstract
Sora is the first large-scale generalist video generation model that garnered significant attention across society. Since its launch by OpenAI in February 2024, no other video generation models have paralleled Sora’s performance or its capacity to support a broad spectrum of video generation tasks. Additionally, there are only a few fully published video generation models, with the majority being closed-source. To address this gap, this paper proposes a new multi-agent framework Mora, which incorporates several advanced visual AI agents to replicate generalist video generation demonstrated by Sora. In particular, Mora can utilize multiple visual agents and successfully mimic Sora’s video generation capabilities in various tasks, such as (1) text-to-video generation, (2) text-conditional image-to-video generation, (3) extend generated videos, (4) video-to-video editing, (5) connect videos and (6) simulate digital worlds. Our extensive experimental results show that Mora achieves performance that is proximate to that of Sora in various tasks. However, there exists an obvious performance gap between our work and Sora when assessed holistically. In summary, we hope this project can guide the future trajectory of video generation through collaborative AI agents.
Figure 1: Samples for text-to-video generation of Mora. Our approach can generate high-resolution, temporally consistent videos from text prompts. The samples shown are 1024$\times$576 resolution over 12 seconds duration at 75 frames in total.
1 Introduction
Since the release of ChatGPT in November 2022, the advent of generative AI technologies has marked a significant transformation, reshaping interactions and integrating deeply into various facets of daily life and industry OpenAI (2023a, b). In the realm of visual AI, the focus has predominantly been on image generation with models such as Midjourney Midjourney (2023), Stable Diffusion Rombach et al. (2022a), and DALL-E 3 Betker et al. (2023) leading the way. In contrast, video generation has remained unexplored, especially when compared to the progress seen in language and image synthesis technologies. Although recent notable video generation models like Pika pik and Gen-2 Gen have demonstrated their capacity to produce diverse and high-quality videos, they have been limited by their ability to create longer-duration videos exceeding 10 seconds. This scenario has seen a revolutionary shift with the introduction of OpenAI’s Sora OpenAI (2024a), marking a new era in video generation capabilities.
Sora, a pioneering text-to-video generative model introduced by OpenAI in February 2024, distinguishes itself by its ability to convert text prompts into detailed videos, demonstrating remarkable potential in replicating physical world dynamics. This model transcends its predecessors by generating videos up to a minute long, closely aligning with the provided text descriptions OpenAI (2024a). Beyond mere text-to-video generation, Sora excels in a range of video tasks as a generalist model, including editing, connecting, and extending footage in ways previously unachieved. In addition, the generated content is known for its multi-view perspectives and fidelity to user instructions, positioning it uniquely amongst video generation models. As we look to the future, the implications of Sora and similar advanced video generation technologies are poised to make profound contributions across various sectors, including but not limited to filmmaking Zhu et al. (2023); Zhuang et al. (2024), robotics Kapelyukh et al. (2023); Liu et al. (2022), and healthcare Awasthi et al. (2023).
Despite its innovative contributions, Sora’s closed-source nature, similar to most video generation models, presents significant challenges for the academic community. The inaccessibility hinders researchers’ ability to replicate or extend Sora’s capabilities. There is a growing trend of attempts to reverse-engineer Sora, with some studies, such as Liu et al. (2024a), proposing potential techniques that might be employed within Sora, including diffusion transformers and spatial patch strategies Sohl-Dickstein et al. (2015); Ho et al. (2020); Bao et al. (2023); Peebles and Xie (2023). Despite these efforts, achieving the same level of performance and adaptability as Sora proves to be an immense challenge. The lack of access to comparable computational power and extensive training datasets further complicates these endeavors.
Table 1: Task comparison between Sora, Mora and other existing models.
Tasks Example Sora Mora Others
Text-to-video Generation
Girdhar et al. (2023); Wang et al. (2023a); Chen et al. (2024); Ma et al. (2024a)
Text-conditional Image-to-Video Generation
Blattmann et al. (2023); pik ; Gen
Extend Generated Videos
Video-to-Video Editing
Molad et al. (2023); Liew et al. (2023); Ceylan et al. (2023)
Connect Videos
Chen et al. (2023a)
Simulate Digital Worlds
To address the limitations of current video generation models, we explore the potential of multi-agent collaboration Hong et al. (2023); Wu et al. (2023a) in accomplishing generalist video generation tasks. We introduce a multi-agent framework, referred to as Mora, that leverages various advanced large models to enable text-to-video capabilities similar to Sora. Specifically, we decompose video generation into several subtasks, with each subtask assigned to a dedicated agent: (1) enhancing prompts provided by the user, (2) generating an image from an input text prompt, (3) editing or refining images based on the enhanced conditioning provided by the text, (4) generating a video from the generated image, and (5) connecting two videos. By automatically organizing agents to loop and permute through these subtasks, Mora can complete a wide array of video generation tasks through a flexible pipeline, thereby meeting the diverse needs of users.
Intuitively, equipping the model with both a starting image and text simplifies the video generation process, as it primarily needs to extrapolate the future progression of the image. This method stands in contrast to direct end-to-end text-to-video approaches Singer et al. (2023); Ho et al. (2022a); Chen et al. (2023b); Wang et al. (2023a); Chen et al. (2024). Our multi-agent collaboration framework distinctively produces an intermediate image or video during inference, enabling the preservation of the visual diversity, style, and quality inherent in the text-to-image model (see examples in Figure 1). This process even facilitates editing capabilities. By effectively coordinating the efforts of text-to-image, image-to-image, image-to-video and video-to-video agents, Mora can adeptly conduct a broad spectrum of video generation tasks while offering superior editing flexibility and visual fidelity, rivaling the performance of established models like Sora. A detailed comparison of tasks between Mora and Sora is presented in Table 1. This comparison demonstrates that, through the collaboration of multiple agents, Mora is capable of accomplishing the video-related tasks that Sora can undertake. This comparison highlights Mora’s adaptability and proficiency in addressing a multitude of video generation challenges.
To comprehensively assess the efficacy of Mora, we use basic metrics in publicly available video generation benchmark Vbench Huang et al. (2023) and self-defined metrics for six tasks, including text-to-video generation, text-conditional image-to-video generation, extending generated videos, video-to-video editing, connecting videos, and simulating digital worlds. Notably, Mora achieves superior performance in the text-to-video generation task than existing open-sourced models, ranking second only to Sora. In the other tasks, Mora also delivers competitive results, underscoring the versatility and general capabilities of our framework.
We summarize our contributions as follows:
  • In this paper, we introduce Mora, a groundbreaking meta-programming framework crafted to enhance multi-agent collaboration. This framework stands out for its structured yet adaptable system of agents, paired with an intuitive interface for the configuration of components and task pipelines. These features position Mora as a prime instrument for pushing forward the boundaries of generalist video generation tasks.
  • Our research reveals that the quality of video generation can be notably improved by leveraging the automated cooperation of multiple agents, including text-to-image, image-to-image, image-to-video, and video-to-video agents. This collaborative process starts with generating an image from text, followed by using both the generated image and the input text to produce a video. The process concludes with further refinement, extension, connection, and editing of the video.
  • Mora stands out for its exceptional performance across six video-related tasks, surpassing existing open-sourced models. This impressive achievement underlines the effectiveness of Mora, showcasing its potential as a versatile framework tailored for generalist video generation. The comprehensive results not only affirm the capabilities of Mora but also position it as a groundbreaking tool in the realm of video generation, promising significant advancements in how video content is created and utilized.
2 Related Work
2.1 Text-to-Video Generation
Generating videos based on textual descriptions has been long discussed. While early efforts in the field were primarily rooted in GANs Wang et al. (2020); Chu et al. (2020) and VQ-VAE Yan et al. (2021), recent breakthroughs in generative video models, driven by foundational work in transformer-based architectures and diffusion models, have advanced academic research. Auto-regressive transformers are early leveraged in video generation Wu et al. (2022); Hong et al. (2022a); Kondratyuk et al. (2023). These models are designed to generate video sequences in a frame-by-frame manner, predicting each new frame based on the previously generated frames. Parallelly, the adaptation of masked language models He et al. (2021) for visual contexts, as demonstrated by Gupta et al. (2022); Villegas et al. (2022); Yu et al. (2023); Gupta et al. (2023), underscores the versatility of transformers in video generation. The recently-proposed VideoPoet Kondratyuk et al. (2023) leverages an auto-regressive language model and can multitask on a variety of video-centric inputs and outputs.
In another line, large-scale diffusion models Sohl-Dickstein et al. (2015); Ho et al. (2020) show competitive performance in video generation Ho et al. (2022a); Singer et al. (2022); Khachatryan et al. (2023); Wu et al. (2023b); Du et al. (2024). By learning to gradually denoise a sample from a normal distribution, diffusion models Sohl-Dickstein et al. (2015); Ho et al. (2020) implement an iterative refinement process for video synthesis. Initially developed for image generation Rombach et al. (2022b); Podell et al. (2023), they have been adapted and extended to handle the complexities of video data. This adaptation began with extending image generation principles to video Ho et al. (2022b, a); Singer et al. (2022), by using a 3D U-Net structure instead of conventional image diffusion U-Net. In the follow-up, latent diffusion models (LDMs) Rombach et al. (2022a) are integrated into video generation Zhou et al. (2022); Chen et al. (2023b); Wang et al. (2023a); Chen et al. (2024), showcasing enhanced capabilities to capture the nuanced dynamics of video content. For instance, Stable Video Diffusion Blattmann et al. (2023) can conduct multi-view synthesis from a single image while Emu Video Girdhar et al. (2023) uses just two diffusion models to generate higher-resolution videos.
Researchers have delved into the potential of diffusion models for a variety of video manipulation tasks. Notably, Dreamix Molad et al. (2023) and MagicEdit Liew et al. (2023) have been introduced for general video editing, utilizing large-scale video-text datasets. Conversely, other models employ pre-trained models for video editing tasks in a zero-shot manner Ceylan et al. (2023); Couairon et al. (2023); Yang et al. (2023); Chai et al. (2023). SEINE Chen et al. (2023a) is specially designed for generative transition between scenes and video prediction. While these above-mentioned end-to-end models exhibit remarkable proficiency in specific areas, they encounter limitations in broadening their capabilities to encompass a wider range of video tasks, particularly those requiring varied types of inputs and the generation of longer-duration videos. The introduction of diffusion transformers Peebles and Xie (2023); Bao et al. (2023); Ma et al. (2024b) further revolutionized video generation, culminating in advanced solutions like Latte Ma et al. (2024a) and Sora OpenAI (2024a). Sora’s ability to produce minute-long videos of high visual quality that faithfully follow human instructions heralds a new era in video generation, promising unprecedented opportunities for creativity and expression in digital media.
2.2 AI Agents
Large models have enabled agents to excel across a broad spectrum of applications, showcasing their versatility and effectiveness. They have greatly advanced collaborative multi-agent structures for multimodal tasks in areas such as scientific research Tang et al. (2024), software development Hong et al. (2023); Qian et al. (2023) and society simulation Park et al. (2023). Compared to individual agents, the collaboration of multiple autonomous agents, each equipped with unique strategies and behaviors and engaged in communication with one another, can tackle more dynamic and complex tasks Guo et al. (2024). Through a cooperative agent framework known as role-playing, CAMEL Li et al. (2024) enables agents to collaborate and solve complex tasks effectively. Park et al. Park et al. (2023) designed a community of 25 generative agents capable of planning, communicating, and forming connections. Liang et al. Liang et al. (2023) have explored the use of multi-agent debates for translation and arithmetic problems, encouraging divergent thinking in large language models. Hong et al. Hong et al. (2023) introduced MetaGPT, which utilizes an assembly line paradigm to assign diverse roles to various agents. In this way, complex tasks can be broken down into subtasks, which makes it easy for many agents working together to complete. Xu et al. Xu et al. (2023) used a multi-agent collaboration strategy to simulate the academic peer review process. AutoGen Wu et al. (2023a) is a generic programming framework which can be used to implement diverse multi-agent applications across different domains, using a variety of agents and conversation patterns.
Motivated by existing works, we extend the principle of collaborative agents to complete vision tasks. By integrating and coordinating agents of various roles such as text-to-image and image-to-video agents, we manage to accomplish multiple video-related tasks in a modular and extensible approach.
3 Mora: A Multi-Agent Framework for Video Generation
As illustrated in Figure 2, Mora aims to provide a framework leveraging advanced AI agents to realize Text-to-Video generation. Sec 3.1 provides the agent definitions and example pipelines enabled by this framework. Sec 3.2 introduces the implementation detail of each agent.
3.1 Agent-based Video Generation
Definition and Specialization of Agents.   The definition of agents enables flexibility in the breakdown of complex work into smaller and more specific tasks. Solving different video generation tasks often requires the collaboration of agents with diverse abilities, each contributing specialized outputs. In our framework, we have 5 basic roles: prompt selection and generation agent, text-to-image generation agent, image-to-image generation agent, and image-to-video generation agent and video-to-video agent.
Figure 2: Illustration of how to use Mora to conduct video-related tasks.
  • Prompt Selection and Generation Agent: Prior to the commencement of the initial image generation, textual prompts undergo a rigorous processing and optimization phase. This critical agent can employ large language models like GPT-4, Llama Achiam et al. (2023); Touvron et al. (2023). It is designed to meticulously analyze the text, extracting pivotal information and actions delineated within, thereby significantly enhancing the relevance and quality of the resultant images. This step ensures that the textual descriptions are thoroughly prepared for an efficient and effective translation into visual representations.
  • Text-to-Image Generation Agent: The text-to-image model Rombach et al. (2022b); Podell et al. (2023) stands at the forefront of translating these enriched textual descriptions into high-quality initial images. Its core functionality revolves around a deep understanding and visualization of complex textual inputs, enabling it to craft detailed and accurate visual counterparts to the provided textual descriptions.
  • Image-to-Image Generation Agent: This agent Brooks et al. (2023) works to modify a given source image in response to specific textual instructions. The core of its functionality lies in its ability to interpret detailed textual prompts with high accuracy and subsequently apply these insights to manipulate the source image accordingly. This involves a detailed recognition of the text’s intent, translating these instructions into visual modifications that can range from subtle alterations to transformative changes. The agent leverages a pre-trained model to bridge the gap between textual description and visual representation, enabling seamless integration of new elements, adjustment of visual styles, or alteration of compositional aspects within the image.
  • Image-to-Video Generation Agent  Blattmann et al. (2023): Following the creation of the initial image, the Video Generation Model is responsible for transitioning the static frame into a vibrant video sequence. This component delves into the analysis of both the content and style of the initial image, serving as the foundation for generating subsequent frames. These frames are meticulously crafted to ensure a seamless narrative flow, resulting in a coherent video that upholds temporal stability and visual consistency throughout. This process highlights the model’s capability to not only understand and replicate the initial image but also to anticipate and execute logical progressions in the scene.
  • Video Connection Agent: Utilizing the Video-to-Video Agent, we create seamless transition videos based on two input videos provided by users. This advanced agent selectively leverages key frames from each input video to ensure a smooth and visually consistent transition between them. It is designed with the capability to accurately identify the common elements and styles across the two videos, thus ensuring a coherent and visually appealing output. This method not only improves the seamless flow between different video segments but also retains the distinct styles of each segment.
Every agent is responsible for the specific input and output. These results can be utilized for different designed tasks.
Approaches.   By setting the agents’ roles and operational skills, we can define basic workflows for different tasks. We design six text-to-video generation tasks: (1) Text-to-video generation, (2) Text-conditional image-to-video generation, (3) Extend generated videos, (4) Video-to-video editing, (5) Connect Videos, and (6) Simulate digital worlds, which are described below.
  • Task 1: Text-to-Video Generation: This task harnesses a detailed textual prompt from the user as the foundation for video creation. The prompt must meticulously detail the envisioned scene. Utilizing this prompt, the Text-to-Image agent utilizes the text, distilling themes and visual details to craft an initial frame. Building upon this foundation, the Image-to-Video component methodically generates a sequence of images. This sequence dynamically evolves to embody the prompt’s described actions or scenes, and each video is derived from the last frame from the previous video, thereby achieving a seamless transition throughout the video.
  • Task 2: Text-conditional Image-to-Video Generation: Task 2 mirrors the operational pipeline of Task 1, with a key distinction. Unlike Task 1 with only texts as inputs, Task 2 integrates both a textual prompt and an initial image into the Text-to-Image agent’s input. This dual-input approach enriches the content generation process, enabling a more nuanced interpretation of the user’s vision.
  • Task 3: Extend Generated Videos: This task focuses on extending the narrative of an existing video sequence. By taking the last frame of an input video as the starting point, the video generation agent crafts a series of new, coherent frames that continue the story. This approach allows for the seamless expansion of video content, creating longer narratives that maintain the consistency and flow of the original sequence.
  • Task 4: Video-to-Video Editing: Task 4 introduces a sophisticated editing capability, leveraging both the Image-to-Image and Image-to-Video agents. The process begins with the Image-to-Image agent, which takes the first frame of an input video and applies edits based on the user’s prompt, achieving the desired modifications. This edited frame then serves as the initial image for the Image-to-Video agent, which generates a new video sequence that reflects the requested obvious or subtle changes, offering a powerful tool for dynamic video editing.
  • Task 5: Connect Videos: The Image-to-Video agent leverages the final frame of the first input video and the initial frame of the second input video to create a seamless transition, producing a new video that smoothly connects the two original videos.
  • Task 6: Simulating Digital Worlds: This task specializes in the whole style changing for video sequences set in digitally styled worlds. By appending the phrase "In digital world style" to the edit prompt, the user instructs the Image-to-Video agent to craft a sequence that embodies the aesthetics and dynamics of a digital realm or utilize the Image-to-Image agent to transfer the real image to digital style. This task pushes the boundaries of video generation, enabling the creation of immersive digital environments that offer a unique visual experience.
3.2 Implementation Detail of Agents
Prompt Selection and Generation.  Currently, GPT-4  Achiam et al. (2023) stands as the most advanced generative model available. By harnessing the capabilities of GPT-4, we are able to generate and meticulously select high-quality prompts. These prompts are detailed and rich in information, facilitating the Text-to-Image generation process by providing the agent with comprehensive guidance.
Text-to-Image Generation.  We utilize the pretrained large text-to-image model to generate a high-quality and representative first image.
The Stable Diffusion XL (SDXL) Podell et al. (2023) is utilized for the first implementation. It introduces a significant evolution in the architecture and methodology of latent diffusion models Rombach et al. (2022b); Meng et al. (2022) for text-to-image synthesis, setting a new benchmark in the field. At the core of its architecture is an enlarged UNet backbone Ronneberger et al. (2015) that is three times larger than those used in previous versions of Stable Diffusion 2 Rombach et al. (2022b). This expansion is principally achieved through an increased number of attention blocks and a broader cross-attention context, facilitated by integrating a dual text encoder system. The first encoder is based on OpenCLIP Ilharco et al. (2021) ViT-bigG Cherti et al. (2023); Radford et al. (2021); Schuhmann et al. (2022), while the second utilizes CLIP ViT-L, allowing for a richer, more nuanced interpretation of textual inputs by concatenating the outputs of these encoders. This architectural innovation is complemented by the introduction of several novel conditioning schemes that do not require external supervision, enhancing the model’s flexibility and capability to generate images across multiple aspect ratios. Moreover, SDXL features a refinement model that employs a post-hoc image-to-image transformation to elevate the visual quality of the generated images. This refinement process utilizes a noising-denoising technique, further polishing the output images without compromising the efficiency or speed of the generation process.
Image-to-Image Generation.  Our initial framework realize uitilze InstructPix2Pix as Image-to-Image generation agent. InstructPix2Pix  Brooks et al. (2023) are intricately designed to enable effective image editing from natural language instructions. At its core, the system integrates the expansive knowledge of two pre-trained models: GPT-3  Brown et al. (2020) for generating editing instructions and edited captions from textual descriptions, and Stable Diffusion  Rombach et al. (2022a) for transforming these text-based inputs into visual outputs. This ingenious approach begins with fine-tuning GPT-3 on a curated dataset of image captions and corresponding edit instructions, resulting in a model that can creatively suggest plausible edits and generate modified captions. Following this, the Stable Diffusion model, augmented with the Prompt-to-Prompt technique, generates pairs of images (before and after the edit) based on the captions produced by GPT-3. The conditional diffusion model at the heart of InstructPix2Pix is then trained on this generated dataset. InstructPix2Pix directly utilizes the text instructions and input image to perform the edit in a single forward pass. This efficiency is further enhanced by employing classifier-free guidance for both the image and instruction conditionings, allowing the model to balance fidelity to the original image with adherence to the editing instructions.
Image-to-Video Generation.  In the Text-to-Video generation agent, video generation agents play an important role in ensuring video quality and consistency.
Our first implementation utilizes the state-of-the-art video generation model Stable Video Diffusion to generate video. The Stable Video Diffusion (SVD)  Blattmann et al. (2023) architecture introduces a cutting-edge approach to generating high-resolution videos by leveraging the strengths of LDMs Stable Diffusion v2.1  Rombach et al. (2022a), originally developed for image synthesis, and extending their capabilities to handle the temporal complexities inherent in video content. At its core, the SVD model follows a three-stage training regime that begins with text-to-image pertaining, where the model learns robust visual representations from a diverse set of images. This foundation allows the model to understand and generate complex visual patterns and textures. In the second stage, video pretraining, the model is exposed to large amounts of video data, enabling it to learn temporal dynamics and motion patterns by incorporating temporal convolution and attention layers alongside its spatial counterparts. This training is conducted on a systematically curated dataset, ensuring the model learns from high-quality and relevant video content. The final stage, high-quality video finetuning, focuses on refining the model’s ability to generate videos with increased resolution and fidelity, using a smaller but higher-quality dataset. This hierarchical training strategy, complemented by a novel data curation process, allows SVD to excel in producing state-of-the-art text-to-video and image-to-video synthesis with remarkable detail, realism, and coherence over time.
Connect Videos.  For the video connection task, we utilize SEINE Chen et al. (2023a) to connect videos. SEINE is constructed upon a pre-trained diffusion-based T2V model, LaVie Wang et al. (2023a) agent. SEINE centered around a random-mask video diffusion model that generates transitions based on textual descriptions. By integrating images of different scenes with text-based control, SEINE produces transition videos that maintain coherence and visual quality. Additionally, the model can be extended for tasks such as image-to-video animation and autoregressive video prediction.
4 Experiments
4.1 Setup
Baseline.  In the text-to-video generation, existing open-sourced works that show competitive performance are employed as the baseline models, including Videocrafter1 Chen et al. (2023c), Show-1 Zhang et al. (2023a), Pika pik , Gen-2 Gen , ModelScope Mod , LaVie-Interpolation, LaVie Wang et al. (2023b) and CogVideo Hong et al. (2022b). In the other five tasks, we compare Mora with Sora.
Basic Metrics.  For text-to-video generation, we employed several metrics from Vbench Huang et al. (2023) for evaluation from two aspects: video quality and video condition consistency.
For video quality measurement, we use six metrics. ❶ Object Consistency, computed by the DINO Caron et al. (2021) feature similarity across frames to assess whether object appearance remains consistent throughout the whole video; ❷ Background Consistency, calculated by CLIP Radford et al. (2021) feature similarity across frames; ❸ Motion Smoothness, which utilizes the motion priors in the video frame interpolation model AMT Li et al. (2023) to evaluate the smoothness of generated motions; ❹ Aesthetic Score, obtained by using the LAION aesthetic predictor lio on each video frame to evaluate the artistic and beauty value perceived by humans, ❺ Dynamic Degree, computed by employing RAFT Teed and Deng (2020) to estimate the degree of dynamics in synthesized videos; ❻ Imaging Quality, calculated by using MUSIQ Ke et al. (2021) image quality predictor trained on SPAQ Fang et al. (2020) dataset.
For measuring video condition consistency, we use two metrics. ❶ Temporal Style, which is determined by utilizing ViCLIP Wang et al. (2023c) to compute the similarity between video features and temporal style description features, thereby reflecting the consistency of the temporal style; ❷ Appearance Style, by calculating the feature similarity between synthesized frames and the input prompt using CLIP Radford et al. (2021), to gauge the consistency of appearance style.
Self-defined Metrics.  For evaluating other tasks, we also define four metrics.
❶ Video-Text Integration $VideoTI$, devised to enhance the quantitative evaluation of the model’s fidelity to textual instructions. It employs LLaVA Liu et al. (2024b) to transfer input image into textual descriptors $T_{i}$ and Video-Llama Zhang et al. (2023b) to transfer videos generated by the model into textual $T_{v}$. The textual representation of the image is prepended with the original instructional text, forming an augmented textual input $T_{mix}$. Both the newly formed text and the video-generated text will be input to BERT Devlin et al. (2018). The embeddings obtained are analyzed for semantic similarity through the computation of cosine similarity, providing a quantitative measurement of the model’s adherence to the given instructions and image.
$VideoTI=cosine(embed_{mix},embed_{v})$ (1)
where $embed_{mix}$ represents the embedding for $T_{mix}$ and $embed_{v}$ for $T_{v}$.
❷ Temporal Consistency $TCON$, designed to measure the coherence between an original video and its extended version, provides a vital tool for assessing the integrity of extended video content. For each input-output video pair, we employ ViCLIP Wang et al. (2023c) video encoder to extract their feature vectors. We then compute cosine similarity to get the score.
$TCON=cosine(V_{input},V_{output})$ (2)
❸ Temporal coherence $Tmean$, by quantifying the correlation between the intermediate generated video and the input videos based on $TCON$.
$Tmean=(TCON_{front}+TCON_{beh})/2$ (3)
where $TCON_{front}$ measures the correlation between the intermediate video and the preceding video in the time series while $TCON_{beh}$ assesses the correlation with the subsequent video. The average of these scores provides an aggregate measure of temporal coherence across the video sequence.
❹ Video Length, to evaluate the models’ efficiency in producing video content, specifically focusing on the maximum duration, measured in seconds.
Table 2: Comparative analysis of text-to-video generation performance between Mora and various other models. The Others category scores are derived from the Hugging Face leaderboard. For Our Mora, the evaluation is based on prompts generated by GPT-4, categorized into three types based on the number of moving objects in the videos: Type I (single object in motion), Type II (two to three objects in motion), and Type III (more than three objects in motion). Differences in the input prompt may account for the superiority of Mora inType II prompt’s scores over those of Sora in the relevant evaluations.
Model Video Quality Object Consistency Background Consistency Motion Smoothness Aesthetic Quality Dynamic Dgree Imaging Quality Temporal Style Video Length(s)
Others
Sora 0.797 0.95 0.96 1.00 0.60 0.69 0.58 0.35 60
VideoCrafter1 0.778 0.95 0.98 0.95 0.63 0.55 0.61 0.26 2
ModelScope 0.758 0.89 0.95 0.95 0.52 0.66 0.58 0.25 2
Show-1 0.751 0.95 0.98 0.98 0.57 0.44 0.59 0.25 3
Pika 0.741 0.96 0.96 0.99 0.63 0.37 0.54 0.24 3
LaVie-Interpolation 0.741 0.92 0.97 0.97 0.54 0.46 0.59 0.26 10
Gen-2 0.733 0.97 0.97 0.99 0.66 0.18 0.63 0.24 4
LaVie 0.746 0.91 0.97 0.96 0.54 0.49 0.61 0.26 3
CogVideo 0.673 0.92 0.95 0.96 0.38 0.42 0.41 0.07 4
Our Mora
Type I 0.782 0.96 0.97 0.99 0.60 0.60 0.57 0.26 12
Type II 0.810 0.94 0.95 0.99 0.57 0.80 0.61 0.26 12
Type III 0.795 0.94 0.93 0.99 0.55 0.80 0.56 0.26 12
Mora 0.792 0.95 0.95 0.99 0.57 0.70 0.59 0.26 12
Implementation Details.  For text-to-video generation, we follow the style of text prompts provided in the official Sora technical report OpenAI (2024b). Subsequently, we employ GPT-4 OpenAI (2023b) to produce more text under a few-shot setting. GPT-4 is also utilized to generate the same number of texts in a zero-shot setting. All generated text prompts are then input into text-to-video models for generating videos. For comparison with Sora, we utilize videos featured on its official website and technical report.
All experiments are conducted on two TESLA A100 GPUs, equipped with a substantial 2$\times$80GB of VRAM. The central processing was handled by 4×AMD EPYC 7552 48-Core Processors. Memory allocation was set at 320GB. The software environment was standardized on PyTorch version 2.0.2 and CUDA 12.2 for video generation and PyTorch version 1.10.2 and CUDA 11.6 for video evaluation.
4.2 Results
Text-to-Video Generation.  The quantitative results are detailed in Table 2. Mora showcases commendable performance across all metrics, making it highly comparable to the top-performing model, Sora, and surpassing the capabilities of other competitors. Specifically, Mora achieved a Video Quality score of 0.792, which closely follows Sora’s leading score of 0.797 and surpasses the current best open-source model like VideoCrafter1. In terms of Object Consistency, Mora scored 0.95, equaling Sora and demonstrating superior consistency in maintaining object identities throughout the videos. For Background Consistency and Motion Smoothness, Mora achieved scores of 0.95 and 0.99, respectively, indicating high fidelity in background stability and fluidity of motion within generated videos. Although Sora achieved 0.96 slightly outperforms Mora in Background Consistency, the margin is minimal. The Aesthetic Quality metric, which assesses the overall visual appeal of the videos, saw Mora scoring 0.57. This score, while not the highest, reflects a competitive stance against other models, with Sora scoring slightly higher at 0.60. Nevertheless, Mora’s performance in Dynamic Degree and Imaging Quality, with scores of 0.70 and 0.59, showcases its strength in generating dynamic, visually compelling content that surpasses all other models. As for Temporal Style, Mora scored 0.26, indicating its robust capability in addressing the temporal aspects of video generation. Although this performance signifies a commendable proficiency, it also highlights a considerable gap between our model and Sora, the leader in this category with a score of 0.35.
In Figure 1, the visual fidelity of Mora’s text-to-video generation is compelling, manifesting high-resolution imagery with acute attention to detail as articulated in the accompanying textual descriptions. The vivid portrayal of scenes, from the liftoff of a rocket to the dynamic coral ecosystem and the urban skateboarding vignette, underscores the system’s adeptness in capturing and translating the essence of the described activities and environments into visually coherent sequences. Notably, the images exude a temporal consistency that speaks to Mora’s nuanced understanding of narrative progression, an essential quality in video synthesis from textual prompts.
Text-conditional Image-to-Video Generation.  Analyzing Table 3, we discern a notable demonstration of Mora’s capabilities in text-conditional image-to-video generation, closely trailing behind Sora. Sora leads with a VideoTI score of 0.90 and Motion Smoothness of 0.99, underscoring its refined alignment of video with the provided text and the fluidity of motion within its generated videos. Mora’s VideoTI score of 0.88 and Motion Smoothness of 0.97, while marginally lower, reflect its robust potential in the accurate interpretation of textual prompts and the generation of smooth motion sequences. Both models exhibit a matched score in Dynamic Degree at 0.75, indicating a comparable proficiency in animating videos with a sense of activity and movement. The Imaging Quality scores show Sora with a slight lead at 0.63 against Mora’s 0.60, suggesting Sora’s subtle superiority in image rendering within video sequences.
Table 3: Text-condtional image-to-video generation. This task evaluates the model’s ability to convert static images and accompanying textual instructions into coherent video sequences. We used 4 pairs of image and text inputs sourced from the official Sora technical report OpenAI (2024b).
Model VideoTI Motion Smoothness Dynamic Dgree Imaging Quality
Sora 0.90 0.99 0.75 0.63
Mora 0.88 0.97 0.75 0.60
Figure 3: Samples for text-conditional image-to-video generation of Mora and Sora. Prompt for the first line image is: Monster Illustration in flat design style of a diverse family of monsters. The group includes a furry brown monster, a sleek black monster with antennas, a spotted green monster, and a tiny polka-dotted monster, all interacting in a playful environment. The second image’s prompt is: An image of a realistic cloud that spells "SORA".
In Figure 3, a qualitative comparison between the video outputs from Sora and Mora reveals that both models adeptly incorporate elements from the input prompt and image. The monster illustration and the cloud spelling "SORA" are well-preserved and dynamically translated into video by both models. Despite quantitative differences, the qualitative results of Mora nearly rival those of Sora, with both models are able to animate the static imagery and narrative elements of the text descriptions into coherent video. This qualitative observation attests to Mora’s capacity to generate videos that closely parallel Sora’s output, achieving a high level of performance in rendering text-conditional imagery into video format while maintaining the thematic and aesthetic essence of the original inputs.
Extend Generated Videos.  The quantitative results from Table 4 reveal that while Sora holds a slight edge over Mora in TCON and Imaging Quality, indicating a higher consistency and fidelity in extending video sequences, Mora’s performance remains close, with a marginal difference of 0.05 in TCON and 0.04 in Imaging Quality. The closeness in Temporal Style scores, 0.24 for Sora and 0.22 for Mora, further signifies that Mora nearly matches Sora in maintaining stylistic continuity over time. Despite Sora’s lead, Mora’s capabilities, particularly in following the temporal style and extending existing videos without significant quality loss, demonstrate its effectiveness in the video extension domain.
Table 4: Extend generated videos. This experiment compares the continuity and quality of the extended video with the same input video given to Sora and Mora respectively. The input video is taken from the full video of the offical Sora technical report OpenAI (2024b), and we take 2 clips as the input video for this task.
Model TCON Imaging Quality Temporal Style
Sora 0.99 0.43 0.24
Mora 0.94 0.39 0.22
Figure 4: Samples for Extend generated video of Mora and Sora.
From a qualitative standpoint, Figure 4 illustrates the competencies of Mora in extending video sequences. Both Sora and Mora adeptly maintain the narrative flow and visual continuity from the original to the extended video. Despite the slight numerical differences highlighted in the quantitative analysis, the qualitative outputs suggest that Mora’s extended videos preserve the essence of the original content with high fidelity. The preservation of dynamic elements such as the rider’s motion and the surrounding environment’s blur effect in the Mora generated sequences showcases its capacity to produce extended videos that are not only coherent but also retain the original’s motion and energy characteristics. This visual assessment underscores Mora’s proficiency in generating extended video content that closely mirrors the original, maintaining the narrative context and visual integrity, thus providing near parity with Sora’s performance.
Table 5: Video-to-video editing. This task examines the model’s ability to edit and transform existing video content in accordance with textual instructions while maintaining visual and stylistic coherence. Data includes 1 input video and 12 instructions from official Sora technical report.
Model Imaging Quality Temporal Style
Sora 0.52 0.24
Mora 0.38 0.23
Video-to-Video Editing.  Table 5 exhibits a comparative analysis of video-to-video editing capabilities between Sora and Mora. Sora secures a higher score in Imaging Quality at 0.52, which suggests a superior capability in preserving the visual details and overall image fidelity during the video editing process. Mora demonstrates capability in video-to-video editing with an Imaging Quality score of 0.38. Although this score reveals a noticeable discrepancy when compare to Sora, it offers valuable insights on pinpointing areas for targeted enhancements in future iterations of the Mora. In the aspect of Temporal Style, both models exhibit proximal performance, with Sora marginally leading at 0.24 compared to Mora’s 0.23. This near-parity underscores the capacity of Mora to closely emulate the stylistic temporal consistency achieved by Sora, ensuring that the edited videos retain a coherent style over time, a crucial aspect of seamless video-to-video editing.
Figure 5: Samples for Video-to-video editing
Table 6: Connect Videos. This task assess the capability of models to seamlessly integrate distinct video clips into a cohesive sequence. $*{}^{*}$ora data we utilize is 5 distinct video-video pairs from Sora technical report website and Mora data we utilize 5 video-video pairs of similar styles.
Model Imaging Quality Tmean
$*{}^{*}$ 0.52 0.64
Mora $*{}^{*}$ 0.42 0.45
Figure 6: Samples for Connect Videos
Upon qualitative evaluation, Figure 5 presents samples from video-to-video editing tasks, wherein both Sora and Mora were instructed to modify the setting to the 1920s style while maintaining the car’s red color. Visually, Sora’s output exhibits a transformation that convincingly alters the modern-day setting into one reminiscent of the 1920s, while carefully preserving the red color of the car. Mora’s transformation, while achieving the task instruction, reveals differences in the execution of the environmental modification, with the sampled frame from generated video suggesting a potential for further enhancement to achieve the visual authenticity displayed by Sora. Nevertheless, Mora ’s adherence to the specified red color of the car underline its ability to follow detailed instructions and enact considerable changes in the video content. This capability, although not as refined as Sora’s, demonstrates Mora’s potential for significant video editing tasks.
Connect Videos.  Quantitative analysis based on the results shown in Table 6 suggests that the model Sora outperforms Mora in terms of Imaging Quality and Tmean. Sora achieves a score of 0.52 in Imaging Quality, indicating its higher fidelity in visual representation compared to Mora’s 0.42. Furthermore, Sora’s superiority is evident in the Temporal coherence aspect with a score of 0.64, which implies that Sora maintains a more consistent visual narrative over time than Mora, which scores 0.45. This quantitative assessment not only solidifies Sora’s position as a superior model in creating high-quality, temporally coherent video sequences but also delineates a trajectory for future enhancements in video connectivity for the Mora framework.
Qualitative analysis based on Figure 6 suggest that, in comparison to Sora’s proficiency in synthesizing intermediate video segments that successfully incorporate background elements from preceding footage and distinct objects from subsequent frames within a single frame, the Mora model demonstrates a blurred background in the intermediate videos, which results in indistinguishable object recognition. Accordingly, this emphasizes the potential for advancing the fidelity of images within the generated intermediate videos as well as enhancing the consistency with the entire video sequence. This would contribute to refining the video connecting process and improving the integration quality of Mora’s model outputs.
Table 7: Simulate digital worlds. This task assesses model’s effectiveness in creating videos that emulate digital or virtual environments, with a focus on preserving the visual details and the distinct appearance style of such worlds. By referring to the video generated by Sora from the official technical report OpenAI (2024b), we used GPT-4 OpenAI (2023b) to build the input prompt for Mora, which contains the "minecraft scene" and the "player character".
Model Imaging Quality Appearance style
Sora 0.62 0.23
Mora 0.52 0.23
Figure 7: Samples for Simulate digital worlds
Simulate Digital Worlds.  In the assessment of digital world simulation capabilities, Table 7 presents a comparative metric analysis between Sora and Mora. Sora exhibits a lead in Imaging Quality with a score of 0.62, indicative of its refined capability to render digital worlds with a higher degree of visual realism and fidelity. Mora, with a score of 0.52, although demonstrating a competent level of performance, falls behind Sora, suggesting areas for improvement in achieving the same level of image clarity and detail. However, both models achieve identical scores in Appearance Style, at 0.23, which reflects a shared ability to adhere to the stylistic parameters of the digital worlds being simulated. This suggests that while there is a difference in the imaging quality, the stylistic translation of textual descriptions into visual aesthetics is accomplished with equivalent proficiency by both models.
Upon qualitative evaluation, Figure 7 presents samples from Simulate digital worlds tasks, wherein both Sora and Mora were instructed to generated video of "Minecraft" scenes. In the top row of frames generated by Sora, we see that the videos maintain high fidelity to the textures and elements typical of digital world aesthetics, characterized by crisp edges, vibrant colors, and clear object definition. The pig and the surrounding environment appear to adhere closely to the style one would expect from a high-resolution game or a digital simulation. These are crucial aspects of performance for Sora, indicating a high-quality synthesis that aligns well with user input while preserving visual consistency and digital authenticity. The bottom row of frames generated by Mora suggests a step towards achieving the digital simulation quality of Sora but with notable differences. Although Mora seems to emulate the digital world’s theme effectively, there is a visible gap in visual fidelity. The images generated by Mora exhibit a slightly muted color palette, less distinct object edges, and a seemingly lower resolution compared to Sora’s output. This suggests that Mora is still in a developmental phase, with its generative capabilities requiring further refinement to reach the performance level of Sora.
5 Discussion
5.1 Strengths of Mora
Innovative Framework and Flexibility.  Mora introduces a groundbreaking multi-agent framework for video generation, significantly advancing the field by enabling a vast array of tasks. This innovative approach not only facilitates text-to-video conversion but also supports the simulation of digital worlds, showcasing unparalleled flexibility and efficiency. Unlike its closed-source counterparts such as Sora, Mora’s open framework design stands out, offering the ability to seamlessly integrate various models. This adaptability allows for the completion of an expanded range of tasks and projects, making Mora an indispensable tool for diverse application scenarios.
Open-Source Contribution.  Mora’s open-source nature is highlighted as a significant contribution to the AI community, encouraging further development and refinement by providing a solid foundation upon which future research can build. This open-source approach not only democratizes access to advanced video generation technologies but also fosters collaboration and innovation in the field. The discussion suggests new ways for future research, including improving the framework’s efficiency, reducing its computational demands during training, and exploring new agent configurations to enhance performance.
5.2 Limitations of Mora
Video Dataset is all you need.  Collecting high-quality video datasets poses significant challenges, primarily due to copyright restrictions on many videos. Unlike images, which can often be easier to gather and utilize for training purposes due to a broader range of available resources and more lenient copyright laws, videos frequently contain copyrighted materials that are not as straightforward to collect or use legally. This issue is compounded by the fact that high-quality videos often come from professional sources, such as movies, television shows, and proprietary game footage, which are rigorously protected by copyright laws. Especially in scenarios involving humans in realistic settings, Mora struggles to generate lifelike movements, such as walking or riding a bicycle. Capturing the subtleties of human motion requires not just any video footage, but high-resolution, smoothly captured sequences that detail every aspect of human kinetics, including the nuances of balance, posture, and interaction with surroundings. Without access to extensive datasets that accurately represent this wide range of human movements, it becomes challenging for video generation models like Mora to replicate these actions convincingly. This limitation underscores the importance of not only the quantity but also the quality and diversity of video datasets in training models to understand and recreate complex human behaviors accurately.
Quality and Length Gaps.  Despite its innovative approach, Mora faces notable challenges, especially when compared to its counterpart, Sora, in terms of video generation quality and capabilities. While Mora is capable of accomplishing tasks similar to those of Sora, the quality of videos generated by Mora falls significantly short, particularly in scenarios involving substantial object movement. Such conditions often introduce a considerable amount of noise, with the quality degradation becoming more pronounced relative to the video length. Moreover, although Mora generally maintains quality for up to 12 seconds of video, any attempt to extend beyond this duration leads to a marked decline in video quality. In contrast, Sora has demonstrated the ability to produce high-quality videos exceeding one minute in length, as showcased in its technical report. This disparity in imaging quality, alongside the limitations in generating videos longer than 12 seconds at a resolution of 1024x576, underscores the urgent need for advancements in Mora’s rendering capabilities and an extension of its video generation parameters.
Instruction Following Capability.  Mora has reached a milestone in its development, now capable of generating videos exceeding 10 seconds from given prompts. However, despite its ability to include all objects specified in the prompts within the generated videos, Mora encounters limitations in executing certain functions. Notably, it struggles with interpreting and rendering the dynamics of motion described in prompts, such as the speed of movement. Additionally, Mora lacks the capability to control the direction of motion—specific directions like left or right for specified objects remain unachievable. This shortfall primarily stems from the system’s foundational approach to video generation, which operates on an Image to Video basis without direct input from textual prompts. Consequently, meeting these specific requirements poses significant challenges, highlighting areas for potential enhancement in Mora’s functionality and interpretative capabilities.
Human Visual Preference Alignment.  The absence of human labeling information within the video domain suggests that experimental results may not always align with human visual preferences, highlighting a significant gap. For example, during connection video task, the generation of transition videos that disintegrate a male figure only to form a female figure represents an exceedingly illogical scenario. This example underscores the necessity for datasets that adhere more closely to physical laws, thereby refining our work and ensuring more realistic and coherent video generation outcomes. Such a framework would enhance the appeal of generated videos to viewers and contribute to establishing comprehensive benchmarks and evaluation criteria for the field.
6 Conclusion
We introduce Mora, a groundbreaking generalist framework for video generation that addresses a range of video-related tasks. Leveraging the collaborative power of multiple agents, Mora marks a considerable advancement in generating videos from textual prompts, establishing new benchmarks for adaptability, efficiency, and output quality in the field of video generation. Our thorough evaluation reveals that Mora not only competes with but also exceeds the capabilities of current leading models in certain areas. However, it has a notable gap with Sora model by OpenAI whose closed-source nature poses considerable challenges for replication and innovation within the academic and professional communities. Our work demonstrates the untapped potential of a meta-programming approach that facilitates intricate collaborations among a variety of agents, each specializing in a segment of the video generation process. In essence, the achievements of Mora not only illustrate the current open source state-of-the-art models in video generation but also illuminate the path forward for the field. As we continue to explore the vast landscape of generative AI, multi-agent collaboration frameworks like Mora will undoubtedly play a pivotal role in unlocking new creative possibilities and applications, from storytelling and content creation to simulation and training. The journey of innovation is far from over, and Mora represents a significant milestone on this ongoing voyage of discovery.
Looking ahead, there are several promising pathways for further research. One such direction involves exploring the integration of more sophisticated natural language understanding capabilities within the agents, potentially allowing for more detailed and context-aware video generations. In addition, the expansion of Mora to incorporate real-time feedback loops could offer interactive video creation experiences, where user inputs could guide the generation process in more dynamic and responsive ways. Furthermore, the challenge of accessibility and computational resource requirements remains a critical barrier to wider adoption and innovation. Future iterations of Mora could benefit from optimizations that reduce these requirements, making advanced video generation technologies more accessible to a broader range of users and developers. In parallel, efforts to create more open and collaborative research environments could accelerate progress in this domain, enabling the community to build upon the foundation laid by Mora framework and other pioneering works.

### v3

@@
 Abstract
-Sora is the first large-scale generalist video generation model that garnered significant attention across society.
-Since its launch by OpenAI in February 2024, no other video generation models have paralleled Sora’s performance or its capacity to support a broad spectrum of video generation tasks.
-Additionally, there are only a few fully published video generation models, with the majority being closed-source.
-To address this gap, this paper proposes a new multi-agent framework Mora, which incorporates several advanced visual AI agents to replicate generalist video generation demonstrated by Sora.
-In particular, Mora can utilize multiple visual agents and successfully mimic Sora’s video generation capabilities in various tasks, such as (1) text-to-video generation, (2) text-conditional image-to-video generation, (3) extend generated videos, (4) video-to-video editing, (5) connect videos and (6) simulate digital worlds.
-Our extensive experimental results show that Mora achieves performance that is proximate to that of Sora in various tasks.
-However, there exists an obvious performance gap between our work and Sora when assessed holistically.
-In summary, we hope this project can guide the future trajectory of video generation through collaborative AI agents.
-Figure 1: Samples for text-to-video generation of Mora.
-Our approach can generate high-resolution, temporally consistent videos from text prompts.
-The samples shown are 1024$\times$576 resolution over 12 seconds duration at 75 frames in total.
+Text-to-video generation has made significant strides, but replicating the capabilities of advanced systems like OpenAI’s Sora remains challenging due to their closed-source nature.
+Existing open-source methods struggle to achieve comparable performance, often hindered by ineffective agent collaboration and inadequate training data quality.
+In this paper, we introduce Mora, a novel multi-agent framework that leverages existing open-source modules to replicate Sora’s functionalities.
+We address these fundamental limitations by proposing three key techniques: (1) multi-agent fine-tuning with a self-modulation factor to enhance inter-agent coordination, (2) a data-free training strategy that uses large models to synthesize training data, and (3) a human-in-the-loop mechanism combined with multimodal large language models for data filtering to ensure high-quality training datasets.
+Our comprehensive experiments on six video generation tasks demonstrate that Mora achieves performance comparable to Sora on VBench Huang et al.
+(2024), outperforming existing open-source methods across various tasks.
+Specifically, in the text-to-video generation task, Mora achieved a Video Quality score of 0.800, surpassing Sora’s 0.797 and outperforming all other baseline models across six key metrics.
+Additionally, in the image-to-video generation task, Mora achieved a perfect Dynamic Degree score of 1.00, demonstrating exceptional capability in enhancing motion realism and achieving higher Imaging Quality than Sora.
+These results highlight the potential of collaborative multi-agent systems and human-in-the-loop mechanisms in advancing text-to-video generation.
+Our code is available at https://github.com/lichao-sun/Mora.
+11footnotetext: Equal contribution
 1 Introduction
-Since the release of ChatGPT in November 2022, the advent of generative AI technologies has marked a significant transformation, reshaping interactions and integrating deeply into various facets of daily life and industry OpenAI (2023a, b).
-In the realm of visual AI, the focus has predominantly been on image generation with models such as Midjourney Midjourney (2023), Stable Diffusion Rombach et al.
-(2022a), and DALL-E 3 Betker et al.
-(2023) leading the way.
-In contrast, video generation has remained unexplored, especially when compared to the progress seen in language and image synthesis technologies.
-Although recent notable video generation models like Pika pik and Gen-2 Gen have demonstrated their capacity to produce diverse and high-quality videos, they have been limited by their ability to create longer-duration videos exceeding 10 seconds.
-This scenario has seen a revolutionary shift with the introduction of OpenAI’s Sora OpenAI (2024a), marking a new era in video generation capabilities.
-Sora, a pioneering text-to-video generative model introduced by OpenAI in February 2024, distinguishes itself by its ability to convert text prompts into detailed videos, demonstrating remarkable potential in replicating physical world dynamics.
-This model transcends its predecessors by generating videos up to a minute long, closely aligning with the provided text descriptions OpenAI (2024a).
-Beyond mere text-to-video generation, Sora excels in a range of video tasks as a generalist model, including editing, connecting, and extending footage in ways previously unachieved.
-In addition, the generated content is known for its multi-view perspectives and fidelity to user instructions, positioning it uniquely amongst video generation models.
-As we look to the future, the implications of Sora and similar advanced video generation technologies are poised to make profound contributions across various sectors, including but not limited to filmmaking Zhu et al.
-(2023); Zhuang et al.
-(2024), robotics Kapelyukh et al.
-(2023); Liu et al.
-(2022), and healthcare Awasthi et al.
-(2023).
-Despite its innovative contributions, Sora’s closed-source nature, similar to most video generation models, presents significant challenges for the academic community.
-The inaccessibility hinders researchers’ ability to replicate or extend Sora’s capabilities.
-There is a growing trend of attempts to reverse-engineer Sora, with some studies, such as Liu et al.
-(2024a), proposing potential techniques that might be employed within Sora, including diffusion transformers and spatial patch strategies Sohl-Dickstein et al.
-(2015); Ho et al.
-(2020); Bao et al.
-(2023); Peebles and Xie (2023).
-Despite these efforts, achieving the same level of performance and adaptability as Sora proves to be an immense challenge.
-The lack of access to comparable computational power and extensive training datasets further complicates these endeavors.
-Table 1: Task comparison between Sora, Mora and other existing models.
-Tasks Example Sora Mora Others
-Text-to-video Generation
-Girdhar et al. (2023); Wang et al. (2023a); Chen et al. (2024); Ma et al. (2024a)
-Text-conditional Image-to-Video Generation
-Blattmann et al. (2023); pik ; Gen
-Extend Generated Videos
-Video-to-Video Editing
-Molad et al. (2023); Liew et al. (2023); Ceylan et al. (2023)
-Connect Videos
-Chen et al. (2023a)
-Simulate Digital Worlds
-To address the limitations of current video generation models, we explore the potential of multi-agent collaboration Hong et al.
-(2023); Wu et al.
-(2023a) in accomplishing generalist video generation tasks.
-We introduce a multi-agent framework, referred to as Mora, that leverages various advanced large models to enable text-to-video capabilities similar to Sora.
-Specifically, we decompose video generation into several subtasks, with each subtask assigned to a dedicated agent: (1) enhancing prompts provided by the user, (2) generating an image from an input text prompt, (3) editing or refining images based on the enhanced conditioning provided by the text, (4) generating a video from the generated image, and (5) connecting two videos.
-By automatically organizing agents to loop and permute through these subtasks, Mora can complete a wide array of video generation tasks through a flexible pipeline, thereby meeting the diverse needs of users.
-Intuitively, equipping the model with both a starting image and text simplifies the video generation process, as it primarily needs to extrapolate the future progression of the image.
-This method stands in contrast to direct end-to-end text-to-video approaches Singer et al.
-(2023); Ho et al.
-(2022a); Chen et al.
-(2023b); Wang et al.
-(2023a); Chen et al.
-(2024).
-Our multi-agent collaboration framework distinctively produces an intermediate image or video during inference, enabling the preservation of the visual diversity, style, and quality inherent in the text-to-image model (see examples in Figure 1).
-This process even facilitates editing capabilities.
-By effectively coordinating the efforts of text-to-image, image-to-image, image-to-video and video-to-video agents, Mora can adeptly conduct a broad spectrum of video generation tasks while offering superior editing flexibility and visual fidelity, rivaling the performance of established models like Sora.
-A detailed comparison of tasks between Mora and Sora is presented in Table 1.
-This comparison demonstrates that, through the collaboration of multiple agents, Mora is capable of accomplishing the video-related tasks that Sora can undertake.
-This comparison highlights Mora’s adaptability and proficiency in addressing a multitude of video generation challenges.
-To comprehensively assess the efficacy of Mora, we use basic metrics in publicly available video generation benchmark Vbench Huang et al.
-(2023) and self-defined metrics for six tasks, including text-to-video generation, text-conditional image-to-video generation, extending generated videos, video-to-video editing, connecting videos, and simulating digital worlds.
-Notably, Mora achieves superior performance in the text-to-video generation task than existing open-sourced models, ranking second only to Sora.
-In the other tasks, Mora also delivers competitive results, underscoring the versatility and general capabilities of our framework.
-We summarize our contributions as follows:
-  • In this paper, we introduce Mora, a groundbreaking meta-programming framework crafted to enhance multi-agent collaboration.
-This framework stands out for its structured yet adaptable system of agents, paired with an intuitive interface for the configuration of components and task pipelines.
-These features position Mora as a prime instrument for pushing forward the boundaries of generalist video generation tasks.
-  • Our research reveals that the quality of video generation can be notably improved by leveraging the automated cooperation of multiple agents, including text-to-image, image-to-image, image-to-video, and video-to-video agents.
-This collaborative process starts with generating an image from text, followed by using both the generated image and the input text to produce a video.
-The process concludes with further refinement, extension, connection, and editing of the video.
-  • Mora stands out for its exceptional performance across six video-related tasks, surpassing existing open-sourced models.
-This impressive achievement underlines the effectiveness of Mora, showcasing its potential as a versatile framework tailored for generalist video generation.
-The comprehensive results not only affirm the capabilities of Mora but also position it as a groundbreaking tool in the realm of video generation, promising significant advancements in how video content is created and utilized.
+Generative AI technologies have significantly transformed various industries, with substantial advancements particularly notable in visual AI through image generation models like Midjourney (Midjourney, 2023), Stable Diffusion 3 (Esser et al., 2024), and DALL-E 3 (Betker et al., 2023).
+These models have demonstrated remarkable capabilities in generating high-quality images from textual descriptions.
+However, progress in text-to-video generation, especially for videos exceeding 10 seconds, has not kept pace.
+Recent developments such as Pika (pik, ) and Gen-3 (Gen, b) have shown potential but are limited to producing short video clips.
+A major breakthrough occurred with OpenAI’s release of Sora in February 2024—a text-to-video model capable of generating minute-long videos that closely align with textual prompts.
+Sora excels in various video tasks, including editing, extending footage, offering multi-view perspectives, and adhering closely to user instructions (OpenAI, 2024a).
+Despite its impressive capabilities, Sora remains a closed-source system, which poses significant barriers to academic research and development.
+Its black-box nature hinders the community’s ability to study, replicate, and build upon its functionalities, thereby slowing progress in the field.
+Attempts to reverse-engineer Sora are exploring potential techniques like diffusion transformers and spatial patch strategies (Sohl-Dickstein et al., 2015; Ho et al., 2020; Bao et al., 2023; Peebles & Xie, 2023), but a large gap still exists due to the intensive computation required to train everything from scratch with a single model.
+To address these challenges, we propose Mora, a novel multi-agent framework that leverages ideas from standardized operating procedures (SOPs) and employs multiple agents using existing open-source modules to replicate the complex functionalities of Sora.
+While SOPs and multi-agent systems have been utilized in text-based tasks (Hong et al., 2023), applying them to text-to-video generation presents unique challenges.
+Naive multi-agent frameworks (Yuan et al., 2024; Xie et al., 2024) often fail because agents lack effective collaboration mechanisms, leading to suboptimal performance.
+Moreover, existing multi-agent video generation approaches struggle to balance pipeline automation with the need for high-quality training data, which is critical for producing high-fidelity videos.
+Collecting high-quality video data for training is time-consuming and computationally expensive (Chen et al., 2024a), and the scarcity of such data hampers the performance of open-source models.
+Furthermore, the quality of available datasets varies widely, making it challenging to train models that can match the performance of proprietary systems like Sora.
+To overcome these challenges, leveraging human-in-the-loop mechanisms for data filtering becomes essential.
+By integrating human expertise in the data synthesis process, we can ensure that the training data is of high quality, which is crucial for training effective video generation models.
+In this paper, we address these fundamental limitations by introducing several key techniques in Mora.
+First, we develop a multi-agent fine-tuning approach with a novel self-modulation factor that enhances coordination among agents, allowing them to collaborate effectively to achieve common goals.
+Second, we employ a data-free training strategy that uses large models to synthesize training data, reducing reliance on large labeled datasets and enabling efficient model training without extensive data collection efforts.
+Third, we leverage human-in-the-loop mechanisms, in combination with multimodal large language models, for data filtering during the training data synthesis process.
+This approach ensures the quality of the synthesized training data, leading to improved performance of the video generation pipeline.
+We summarize the overall pipeline and supported tasks in Figure 1.
+By integrating human-in-the-loop mechanisms with multimodal large language models for data filtering in our data-free training strategy, we significantly enhance the quality of the training data and, consequently, the generated videos.
+These techniques collectively improve inter-agent and agent-human collaboration, enhance the quality and diversity of the generated videos, and expand the system’s capabilities to match those of Sora.
+Our comprehensive experiments on six video generation tasks demonstrate that Mora achieves performance comparable to Sora on VBench (Huang et al., 2024), closely approaching Sora and outperforming existing open-source methods across various tasks.
+Specifically, in the text-to-video generation, Mora achieves a Video Quality score of 0.800, surpassing Sora’s 0.797, and outperforms all other baselines.
+In the Image-to-Video Generation task, Mora matches Sora in Video-Text Integration (0.90) and Motion Smoothness (0.98 vs.
+0.99), and surpasses Sora in Imaging Quality (0.67 vs.
+0.63) and Dynamic Degree (1.00 vs.
+0.75).
+Additionally, in Video-to-Video Editing, Mora matches Sora in both Imaging Quality and Temporal Style, each scoring 0.52 and 0.24, respectively.
+These results demonstrate Mora’s ability to not only replicate but also enhance the functionalities of Sora, providing a promising platform for future research.
+Figure 1: Illustration of SOPs to conduct video-related tasks in Mora.
+We summarize our main contributions as follows:
+  • We introduce Mora, a novel multi-agent framework that leverages existing open-source modules to replicate the functionalities of Sora in text-to-video generation.
+  • We address fundamental challenges in agent collaboration for video generation through three key designs: (1) a self-modulated multi-agent fine-tuning approach with dynamic modulation factor for enhanced coordination, (2) human-in-the-loop control mechanisms enabling real-time adjustments, and (3) a data-free training strategy using large models to synthesize diverse training data.
+These techniques significantly improve inter-agent cooperation and output quality in multi-agent video generation systems.
+  • We demonstrate through extensive experiments that Mora achieves performance comparable to Sora on standard benchmarks, advancing the field of text-to-video generation and providing an accessible platform for future research.
 2 Related Work
 2.1 Text-to-Video Generation
-Generating videos based on textual descriptions has been long discussed.
-While early efforts in the field were primarily rooted in GANs Wang et al.
-(2020); Chu et al.
-(2020) and VQ-VAE Yan et al.
-(2021), recent breakthroughs in generative video models, driven by foundational work in transformer-based architectures and diffusion models, have advanced academic research.
-Auto-regressive transformers are early leveraged in video generation Wu et al.
-(2022); Hong et al.
-(2022a); Kondratyuk et al.
-(2023).
-These models are designed to generate video sequences in a frame-by-frame manner, predicting each new frame based on the previously generated frames.
-Parallelly, the adaptation of masked language models He et al.
-(2021) for visual contexts, as demonstrated by Gupta et al.
-(2022); Villegas et al.
-(2022); Yu et al.
-(2023); Gupta et al.
-(2023), underscores the versatility of transformers in video generation.
-The recently-proposed VideoPoet Kondratyuk et al.
-(2023) leverages an auto-regressive language model and can multitask on a variety of video-centric inputs and outputs.
-In another line, large-scale diffusion models Sohl-Dickstein et al.
-(2015); Ho et al.
-(2020) show competitive performance in video generation Ho et al.
-(2022a); Singer et al.
-(2022); Khachatryan et al.
-(2023); Wu et al.
-(2023b); Du et al.
-(2024).
-By learning to gradually denoise a sample from a normal distribution, diffusion models Sohl-Dickstein et al.
-(2015); Ho et al.
-(2020) implement an iterative refinement process for video synthesis.
-Initially developed for image generation Rombach et al.
-(2022b); Podell et al.
-(2023), they have been adapted and extended to handle the complexities of video data.
-This adaptation began with extending image generation principles to video Ho et al.
-(2022b, a); Singer et al.
-(2022), by using a 3D U-Net structure instead of conventional image diffusion U-Net.
-In the follow-up, latent diffusion models (LDMs) Rombach et al.
-(2022a) are integrated into video generation Zhou et al.
-(2022); Chen et al.
-(2023b); Wang et al.
-(2023a); Chen et al.
-(2024), showcasing enhanced capabilities to capture the nuanced dynamics of video content.
-For instance, Stable Video Diffusion Blattmann et al.
-(2023) can conduct multi-view synthesis from a single image while Emu Video Girdhar et al.
-(2023) uses just two diffusion models to generate higher-resolution videos.
+Generating videos based on textual descriptions has been a long-discussed topic.
+While early efforts in the field were primarily rooted in GANs (Wang et al., 2020; Chu et al., 2020) and VQ-VAE (Yan et al., 2021), recent breakthroughs in generative video models, driven by foundational work in transformer-based architectures and diffusion models, have significantly advanced academic research.
+Auto-regressive transformers are early leveraged in video generation (Wu et al., 2022; Hong et al., 2022b; Kondratyuk et al., 2023).
+These models are designed to generate video sequences frame by frame, predicting each new frame based on the previously generated frames.
+Parallelly, the adaptation of masked language models (He et al., 2021) for visual contexts, as demonstrated by (Gupta et al., 2022; Villegas et al., 2022; Yu et al., 2023; Gupta et al., 2023), underscores the versatility of transformers in video generation.
+The recently proposed VideoPoet (Kondratyuk et al., 2023) leverages an auto-regressive language model and can multitask on a variety of video-centric inputs and outputs.
+In another line, large-scale diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) show competitive performance in video generation (Ho et al., 2022a; Singer et al., 2022; Khachatryan et al., 2023; Wu et al., 2023a; Du et al., 2024).
+By learning to gradually denoise a sample from a normal distribution, diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) implement an iterative refinement process for video synthesis.
+Initially developed for image generation (Rombach et al., 2022b; Podell et al., 2023), they have been adapted and extended to handle the complexities of video data.
+This adaptation began with extending image generation principles to video (Ho et al., 2022b; a; Singer et al., 2022), by using a 3D U-Net structure instead of conventional image diffusion U-Net.
+In the follow-up, latent diffusion models (LDMs) (Rombach et al., 2022a) are integrated into video generation (Zhou et al., 2022; Chen et al., 2023b; Wang et al., 2023a; Chen et al., 2024a), showcasing enhanced capabilities to capture the nuanced dynamics of video content.
+For instance, Stable Video Diffusion (Blattmann et al., 2023) can conduct multi-view synthesis from a single image while Emu Video (Girdhar et al., 2023) uses just two diffusion models to generate higher-resolution videos.
 Researchers have delved into the potential of diffusion models for a variety of video manipulation tasks.
-Notably, Dreamix Molad et al.
-(2023) and MagicEdit Liew et al.
-(2023) have been introduced for general video editing, utilizing large-scale video-text datasets.
-Conversely, other models employ pre-trained models for video editing tasks in a zero-shot manner Ceylan et al.
-(2023); Couairon et al.
-(2023); Yang et al.
-(2023); Chai et al.
-(2023).
-SEINE Chen et al.
-(2023a) is specially designed for generative transition between scenes and video prediction.
-While these above-mentioned end-to-end models exhibit remarkable proficiency in specific areas, they encounter limitations in broadening their capabilities to encompass a wider range of video tasks, particularly those requiring varied types of inputs and the generation of longer-duration videos.
-The introduction of diffusion transformers Peebles and Xie (2023); Bao et al.
-(2023); Ma et al.
-(2024b) further revolutionized video generation, culminating in advanced solutions like Latte Ma et al.
-(2024a) and Sora OpenAI (2024a).
-Sora’s ability to produce minute-long videos of high visual quality that faithfully follow human instructions heralds a new era in video generation, promising unprecedented opportunities for creativity and expression in digital media.
+Notably, Dreamix (Molad et al., 2023) and MagicEdit (Liew et al., 2023) have been introduced for general video editing, utilizing large-scale video-text datasets.
+Conversely, other models employ pre-trained models for video editing tasks in a zero-shot manner (Ceylan et al., 2023; Couairon et al., 2023; Yang et al., 2023; Chai et al., 2023).
+SEINE (Chen et al., 2023c) is specially designed for generative transition between scenes and video prediction.
+The introduction of diffusion transformers (Peebles & Xie, 2023; Bao et al., 2023; Ma et al., 2024a) further revolutionized video generation, culminating in advanced solutions like Latte (Ma et al., 2024b) and Sora (OpenAI, 2024a).
+There is also utilization in a specific domain such as Bora (Sun et al., 2024) in biomedical scenarios.
+Sora’s ability to produce minute-long videos of high visual quality that faithfully follow human instructions heralds a new era in video generation.
 2.2 AI Agents
 Large models have enabled agents to excel across a broad spectrum of applications, showcasing their versatility and effectiveness.
-They have greatly advanced collaborative multi-agent structures for multimodal tasks in areas such as scientific research Tang et al.
-(2024), software development Hong et al.
-(2023); Qian et al.
-(2023) and society simulation Park et al.
-(2023).
-Compared to individual agents, the collaboration of multiple autonomous agents, each equipped with unique strategies and behaviors and engaged in communication with one another, can tackle more dynamic and complex tasks Guo et al.
-(2024).
-Through a cooperative agent framework known as role-playing, CAMEL Li et al.
-(2024) enables agents to collaborate and solve complex tasks effectively.
-Park et al.
+They have greatly advanced collaborative multi-agent structures for multimodal tasks in areas such as scientific research (Tang et al., 2024), software development (Hong et al., 2023; Qian et al., 2023) and society simulation (Park et al., 2023).
+Compared to individual agents, the collaboration of multiple autonomous agents, each equipped with unique strategies and behaviors and engaged in communication with one another, can tackle more dynamic and complex tasks (Guo et al., 2024).
+Through a cooperative agent framework known as role-playing,  Li et al.
+(2024b) enables agents to collaborate and solve complex tasks effectively.
 Park et al.
@@
 Liang et al.
-Liang et al.
 (2023) have explored the use of multi-agent debates for translation and arithmetic problems, encouraging divergent thinking in large language models.
-Hong et al.
 Hong et al.
@@
 Xu et al.
-Xu et al.
 (2023) used a multi-agent collaboration strategy to simulate the academic peer review process.
-AutoGen Wu et al.
-(2023a) is a generic programming framework which can be used to implement diverse multi-agent applications across different domains, using a variety of agents and conversation patterns.
-Motivated by existing works, we extend the principle of collaborative agents to complete vision tasks.
-By integrating and coordinating agents of various roles such as text-to-image and image-to-video agents, we manage to accomplish multiple video-related tasks in a modular and extensible approach.
+Besides, AutoGen (Wu et al., 2023b) is a generic programming framework that can be used to implement diverse multi-agent applications across different domains, using a variety of agents and conversation patterns.
+This motivates our focus on applying multi-agent frameworks on text-to-video generation tasks, enabling agents to collaborate seamlessly from project inception to completion.
 3 Mora: A Multi-Agent Framework for Video Generation
-As illustrated in Figure 2, Mora aims to provide a framework leveraging advanced AI agents to realize Text-to-Video generation.
-Sec 3.1 provides the agent definitions and example pipelines enabled by this framework.
-Sec 3.2 introduces the implementation detail of each agent.
-3.1 Agent-based Video Generation
+Current text-to-video generation models directly generate videos from textual inputs, which prevents users from supervising key aspects of video quality, style, and other important elements in real-time.
+To address this limitation, we propose a novel multi-agent system coupled with a self-modulated training algorithm specifically designed for the video generation task.
+In the subsequent sections, we outline our approach in detail.
+Section 3.1 describes the problem and design of our multi-agent system and the architecture of our model.
+Finally, in Section 3.2 and Appendix A.7, we present our data-free multi-agent fine-tuning method.
+3.1 Agent Architecture of Mora
+In this section, we introduce the problem and our multi-agent video generation system.
+To address the complexity of generating high-quality, long-duration videos, we propose a multi-agent framework where the generative model $G$ is composed of $n$ collaborating agents $\{A_{1},A_{2},\cdots,A_{n}\}$.
+As shown in Figure 1, each agent specializes in a specific subtask and collaborates together within the video generation pipeline.
+We further introduce the definition of each agent below.
+Problem Definition
+Let $P\in\mathcal{P}$ denote a textual prompt from the space of all possible prompts $\mathcal{P}$, describing the desired video content.
+A video $V$ is represented as a sequence of frames $V=\{F_{1},F_{2},\cdots,F_{T}\}$, where each $F_{t}\in\mathbb{R}^{H\times W\times C}$ corresponds to an image at time step $t$, with height $H$, width $W$, and $C$ color channels.
+Our goal is to generate extended-length videos that are semantically consistent with the textual prompt while exhibiting high visual quality and temporal coherence.
+Formally, we aim to learn a generative model $G:\mathcal{P}\rightarrow\mathcal{V}$ that maps a textual prompt $P$ to a video $V$ in the space of all possible videos $\mathcal{V}$: $V=G(P)$.
+The quality of the generated video can be assessed using a set of metrics =$\mathcal{M}=\{m_{1},m_{2},...,m_{K}\}$, where each $m_{i}:\mathcal{V}\times\mathcal{P}\rightarrow\mathbb{R}$ evaluates a specific aspect of the video (e.g., visual quality, temporal consistency, semantic alignment with the prompt).
+Our objective is to maximize these quality metrics while ensuring diversity in the generated videos.
+In our multi-agent framework, $G$ is decomposed into a set of specialized agents $\{A_{1},A_{2},...,A_{N}\}$, each responsible for a specific subtask in the video generation process.
+These agents collaborate to produce the final video output, allowing for more granular control of the generation process.
 Definition and Specialization of Agents.
-The definition of agents enables flexibility in the breakdown of complex work into smaller and more specific tasks.
-Solving different video generation tasks often requires the collaboration of agents with diverse abilities, each contributing specialized outputs.
-In our framework, we have 5 basic roles: prompt selection and generation agent, text-to-image generation agent, image-to-image generation agent, and image-to-video generation agent and video-to-video agent.
-Figure 2: Illustration of how to use Mora to conduct video-related tasks.
-  • Prompt Selection and Generation Agent: Prior to the commencement of the initial image generation, textual prompts undergo a rigorous processing and optimization phase.
-This critical agent can employ large language models like GPT-4, Llama Achiam et al.
-(2023); Touvron et al.
-(2023).
-It is designed to meticulously analyze the text, extracting pivotal information and actions delineated within, thereby significantly enhancing the relevance and quality of the resultant images.
-This step ensures that the textual descriptions are thoroughly prepared for an efficient and effective translation into visual representations.
-  • Text-to-Image Generation Agent: The text-to-image model Rombach et al.
-(2022b); Podell et al.
-(2023) stands at the forefront of translating these enriched textual descriptions into high-quality initial images.
-Its core functionality revolves around a deep understanding and visualization of complex textual inputs, enabling it to craft detailed and accurate visual counterparts to the provided textual descriptions.
-  • Image-to-Image Generation Agent: This agent Brooks et al.
-(2023) works to modify a given source image in response to specific textual instructions.
-The core of its functionality lies in its ability to interpret detailed textual prompts with high accuracy and subsequently apply these insights to manipulate the source image accordingly.
-This involves a detailed recognition of the text’s intent, translating these instructions into visual modifications that can range from subtle alterations to transformative changes.
-The agent leverages a pre-trained model to bridge the gap between textual description and visual representation, enabling seamless integration of new elements, adjustment of visual styles, or alteration of compositional aspects within the image.
-  • Image-to-Video Generation Agent  Blattmann et al.
-(2023): Following the creation of the initial image, the Video Generation Model is responsible for transitioning the static frame into a vibrant video sequence.
-This component delves into the analysis of both the content and style of the initial image, serving as the foundation for generating subsequent frames.
-These frames are meticulously crafted to ensure a seamless narrative flow, resulting in a coherent video that upholds temporal stability and visual consistency throughout.
-This process highlights the model’s capability to not only understand and replicate the initial image but also to anticipate and execute logical progressions in the scene.
-  • Video Connection Agent: Utilizing the Video-to-Video Agent, we create seamless transition videos based on two input videos provided by users.
-This advanced agent selectively leverages key frames from each input video to ensure a smooth and visually consistent transition between them.
-It is designed with the capability to accurately identify the common elements and styles across the two videos, thus ensuring a coherent and visually appealing output.
-This method not only improves the seamless flow between different video segments but also retains the distinct styles of each segment.
-Every agent is responsible for the specific input and output. These results can be utilized for different designed tasks.
-Approaches.
-By setting the agents’ roles and operational skills, we can define basic workflows for different tasks.
-We design six text-to-video generation tasks: (1) Text-to-video generation, (2) Text-conditional image-to-video generation, (3) Extend generated videos, (4) Video-to-video editing, (5) Connect Videos, and (6) Simulate digital worlds, which are described below.
-  • Task 1: Text-to-Video Generation: This task harnesses a detailed textual prompt from the user as the foundation for video creation.
-The prompt must meticulously detail the envisioned scene.
-Utilizing this prompt, the Text-to-Image agent utilizes the text, distilling themes and visual details to craft an initial frame.
-Building upon this foundation, the Image-to-Video component methodically generates a sequence of images.
-This sequence dynamically evolves to embody the prompt’s described actions or scenes, and each video is derived from the last frame from the previous video, thereby achieving a seamless transition throughout the video.
-  • Task 2: Text-conditional Image-to-Video Generation: Task 2 mirrors the operational pipeline of Task 1, with a key distinction.
-Unlike Task 1 with only texts as inputs, Task 2 integrates both a textual prompt and an initial image into the Text-to-Image agent’s input.
-This dual-input approach enriches the content generation process, enabling a more nuanced interpretation of the user’s vision.
-  • Task 3: Extend Generated Videos: This task focuses on extending the narrative of an existing video sequence.
-By taking the last frame of an input video as the starting point, the video generation agent crafts a series of new, coherent frames that continue the story.
-This approach allows for the seamless expansion of video content, creating longer narratives that maintain the consistency and flow of the original sequence.
-  • Task 4: Video-to-Video Editing: Task 4 introduces a sophisticated editing capability, leveraging both the Image-to-Image and Image-to-Video agents.
-The process begins with the Image-to-Image agent, which takes the first frame of an input video and applies edits based on the user’s prompt, achieving the desired modifications.
-This edited frame then serves as the initial image for the Image-to-Video agent, which generates a new video sequence that reflects the requested obvious or subtle changes, offering a powerful tool for dynamic video editing.
-  • Task 5: Connect Videos: The Image-to-Video agent leverages the final frame of the first input video and the initial frame of the second input video to create a seamless transition, producing a new video that smoothly connects the two original videos.
-  • Task 6: Simulating Digital Worlds: This task specializes in the whole style changing for video sequences set in digitally styled worlds.
-By appending the phrase "In digital world style" to the edit prompt, the user instructs the Image-to-Video agent to craft a sequence that embodies the aesthetics and dynamics of a digital realm or utilize the Image-to-Image agent to transfer the real image to digital style.
-This task pushes the boundaries of video generation, enabling the creation of immersive digital environments that offer a unique visual experience.
-3.2 Implementation Detail of Agents
-Prompt Selection and Generation.
-Currently, GPT-4  Achiam et al.
-(2023) stands as the most advanced generative model available.
-By harnessing the capabilities of GPT-4, we are able to generate and meticulously select high-quality prompts.
-These prompts are detailed and rich in information, facilitating the Text-to-Image generation process by providing the agent with comprehensive guidance.
-Text-to-Image Generation.  We utilize the pretrained large text-to-image model to generate a high-quality and representative first image.
-The Stable Diffusion XL (SDXL) Podell et al.
-(2023) is utilized for the first implementation.
-It introduces a significant evolution in the architecture and methodology of latent diffusion models Rombach et al.
-(2022b); Meng et al.
-(2022) for text-to-image synthesis, setting a new benchmark in the field.
-At the core of its architecture is an enlarged UNet backbone Ronneberger et al.
-(2015) that is three times larger than those used in previous versions of Stable Diffusion 2 Rombach et al.
-(2022b).
-This expansion is principally achieved through an increased number of attention blocks and a broader cross-attention context, facilitated by integrating a dual text encoder system.
-The first encoder is based on OpenCLIP Ilharco et al.
-(2021) ViT-bigG Cherti et al.
-(2023); Radford et al.
-(2021); Schuhmann et al.
-(2022), while the second utilizes CLIP ViT-L, allowing for a richer, more nuanced interpretation of textual inputs by concatenating the outputs of these encoders.
-This architectural innovation is complemented by the introduction of several novel conditioning schemes that do not require external supervision, enhancing the model’s flexibility and capability to generate images across multiple aspect ratios.
-Moreover, SDXL features a refinement model that employs a post-hoc image-to-image transformation to elevate the visual quality of the generated images.
-This refinement process utilizes a noising-denoising technique, further polishing the output images without compromising the efficiency or speed of the generation process.
-Image-to-Image Generation.
-Our initial framework realize uitilze InstructPix2Pix as Image-to-Image generation agent.
-InstructPix2Pix  Brooks et al.
-(2023) are intricately designed to enable effective image editing from natural language instructions.
-At its core, the system integrates the expansive knowledge of two pre-trained models: GPT-3  Brown et al.
-(2020) for generating editing instructions and edited captions from textual descriptions, and Stable Diffusion  Rombach et al.
-(2022a) for transforming these text-based inputs into visual outputs.
-This ingenious approach begins with fine-tuning GPT-3 on a curated dataset of image captions and corresponding edit instructions, resulting in a model that can creatively suggest plausible edits and generate modified captions.
-Following this, the Stable Diffusion model, augmented with the Prompt-to-Prompt technique, generates pairs of images (before and after the edit) based on the captions produced by GPT-3.
-The conditional diffusion model at the heart of InstructPix2Pix is then trained on this generated dataset.
-InstructPix2Pix directly utilizes the text instructions and input image to perform the edit in a single forward pass.
-This efficiency is further enhanced by employing classifier-free guidance for both the image and instruction conditionings, allowing the model to balance fidelity to the original image with adherence to the editing instructions.
-Image-to-Video Generation.  In the Text-to-Video generation agent, video generation agents play an important role in ensuring video quality and consistency.
-Our first implementation utilizes the state-of-the-art video generation model Stable Video Diffusion to generate video.
-The Stable Video Diffusion (SVD)  Blattmann et al.
-(2023) architecture introduces a cutting-edge approach to generating high-resolution videos by leveraging the strengths of LDMs Stable Diffusion v2.1  Rombach et al.
-(2022a), originally developed for image synthesis, and extending their capabilities to handle the temporal complexities inherent in video content.
-At its core, the SVD model follows a three-stage training regime that begins with text-to-image pertaining, where the model learns robust visual representations from a diverse set of images.
-This foundation allows the model to understand and generate complex visual patterns and textures.
-In the second stage, video pretraining, the model is exposed to large amounts of video data, enabling it to learn temporal dynamics and motion patterns by incorporating temporal convolution and attention layers alongside its spatial counterparts.
-This training is conducted on a systematically curated dataset, ensuring the model learns from high-quality and relevant video content.
-The final stage, high-quality video finetuning, focuses on refining the model’s ability to generate videos with increased resolution and fidelity, using a smaller but higher-quality dataset.
-This hierarchical training strategy, complemented by a novel data curation process, allows SVD to excel in producing state-of-the-art text-to-video and image-to-video synthesis with remarkable detail, realism, and coherence over time.
-Connect Videos.
-For the video connection task, we utilize SEINE Chen et al.
-(2023a) to connect videos.
-SEINE is constructed upon a pre-trained diffusion-based T2V model, LaVie Wang et al.
-(2023a) agent.
-SEINE centered around a random-mask video diffusion model that generates transitions based on textual descriptions.
-By integrating images of different scenes with text-based control, SEINE produces transition videos that maintain coherence and visual quality.
-Additionally, the model can be extended for tasks such as image-to-video animation and autoregressive video prediction.
+The specialization of agents enables flexibility in the breakdown of complex work into smaller and more focused tasks, as depicted in Figure 1.
+In our framework, we have five agents: prompt selection and generation agent $A_{1}$, text-to-image generation agent $A_{2}$, image-to-image generation agent $A_{3}$, image-to-video generation agent $A_{4}$, and video-to-video agent $A_{5}$.
+We present brief descriptions below, and detailed definitions can be found in Appendix A.5.
+  • Prompt Selection and Generation Agent ($A_{1}$): This agent employs large language models like GPT-4 and Llama (Achiam et al., 2023; Touvron et al., 2023) to analyze and enhance textual prompts, extracting key information and actions to optimize image relevance and quality.
+  • Text-to-Image Generation Agent ($A_{2}$): Utilizing models such as those by Rombach et al.
+(2022b) and Podell et al.
+(2023), this agent translates enriched textual descriptions into high-quality images by deeply understanding and visualizing complex inputs.
+  • Image-to-Image Generation Agent ($A_{3}$): Referencing Brooks et al.
+(2023), this agent modifies source images based on detailed textual instructions, accurately interpreting and applying these to make visual alterations ranging from subtle to transformative.
+  • Image-to-Video Generation Agent ($A_{4}$): As described by Blattmann et al.
+(2023), this model transitions static images into dynamic video sequences by analyzing content and style, ensuring temporal stability and visual consistency.
+  • Video Connection Agent ($A_{5}$): This agent creates seamless transition videos from two input videos by leveraging keyframes and identifying common elements and styles, ensuring coherent and visually appealing outputs.
+General Structure.
+Mora model structure is depicted in Figure 1.
+Specifically, given a user input $T$, the Prompt Enhancement Agent ($A_{1}$) first refines $T$ into a form better suited for video generation.
+The enhanced prompt is then passed to the Text-to-Image Agent ($A_{2}$) to generate the first frame of the video.
+At this stage, the user can review and confirm whether the tone and quality of the frame meet their expectations.
+If not, the user can either request a re-generation or pass the frame to the Image-to-Image Agent ($A_{3}$) for adjustments.
+Once the desired first frame is finalized, it is forwarded to the Image-to-Video Agent ($A_{4}$) to generate a high-quality video that aligns with the user’s requirements.
+This step-by-step, independently controllable, and interactive process ensures that the final video more closely meets the user’s expectations while maintaining high quality.
+In cases where a user wishes to generate a continuous video from different video clips, $A_{5}$ analyzes the final frame of the preceding video and the initial frame of the next and ensures a smooth blending between them.
+Moreover, this procedural design ensures that the generation process does not have to start from scratch, and the human-in-the-loop technique makes the entire pipeline more controllable, as detailed in Appendix A.4.
+It enables our framework to handle various tasks, such as image-to-video generation, and even video extension and stitching.
+For more details about supported tasks, please refer to Appendix A.3.
+In addition to prompt-based generation, Mora structure also supports task-wise model fine-tuning, ensuring that agents can effectively follow instructions and consistently produce high-quality content.
+3.2 Self-Modulated Multi-Agent Finetuning
+In this section, we introduce our proposed multi-agent finetuning design, based on the previously described model structure.
+Directly prompting each agent does not account for the downward transmission of information, which could lead to inefficiencies or errors in communication between agents.
+Additionally, the impact of each agent on the final result is not uniform.
+To address these issues, we adopt an end-to-end training approach.
+Our proposed training procedure involves (1) a self-modulation factor to enhance inter-agent coordination, (2) a data-free training strategy to synthesize training data, (3) LLM selection with human-in-the-loop to control the training data quality.
+In the following sections, we provide a detailed explanation of each component.
+Self-modulated Fine-tuning Algorithm.
+Previous methods primarily employ direct end-to-end fine-tuning across the entire task procedure, while others may fine-tune individual agents based on specific tasks.
+However, both approaches can influence model performance: (1) end-to-end fine-tuning may result in improper loss allocation for each agent, and (2) fine-tuning agents individually can lead to misaligned distributions.
+To address the limitations of existing fine-tuning approaches, we propose a self-modulated fine-tuning algorithm specifically designed for our multi-agent model structure.
+This method aims to enhance coordination among agents and improve overall performance by introducing a modulation embedding that dynamically adjusts the influence of each agent during the generation process.
+The key motivations behind our approach are to balance the impact of each agent on the final output, improve inter-agent communication and coordination, and allow for dynamic adjustment of agent contributions based on the current task and intermediate outputs.
+Our design introduces a modulation embedding that is concatenated with the text embedding of the enhanced prompt, allowing for fine-grained control over the generation process.
+This embedding is optimized alongside the model parameters during training, enabling the system to learn optimal coordination strategies.
+By doing so, we ensure that each agent can adapt its output based on the state of the preceding agent, leading to more coherent and high-quality video generation.
+The implementation of our self-modulated fine-tuning algorithm begins with the initialization of a modulation embedding $\{\mathbf{z}_{i}\}$ using the text embedding of a special token (Ning et al., 2023): $\{\mathbf{z}_{i}=\mathcal{E}_{i}(\text{"{[Mod]}"})\}$] " ) }, where $\mathcal{E}_{i}$ is the text encoder of agent $i$.
+For each enhanced prompt $P_{\text{enh}}$, we compute its text embedding $\{\mathbf{e}_{i}=\mathcal{E}_{i}(P_{\text{enh}})\}$ and concatenate it with the modulation embedding: $\{\tilde{\mathbf{e}}_{i}=[\mathbf{e}_{i};\mathbf{z}_{i}]\}$.
+This combined embedding serves as input to the agents.
+Each agent $\mathcal{M}_{i}$ in the system processes its input and produces an output $O_{i}$.
+For the first agent, $O_{1}=\mathcal{M}_{1}(\tilde{\mathbf{e}})$, and for subsequent agents ($i>1$), $O_{i}=\mathcal{M}_{i}(O_{i-1},\tilde{\mathbf{e}})$.
+The modulation factor, which influences the impact of each agent, is calculated as the L2 norm of $\mathbf{z}_{i}$: $||\mathbf{z}_{i}||_{2}=\sqrt{\sum_{k}z_{k}^{2}}$.
+During training, we minimize the total loss $L_{\text{total}}=L(O_{n},V_{\text{target}})$ between the final output $O_{n}$ and the target video $V_{\text{target}}$.
+Both the model parameters $\theta_{i}$ of each agent and the modulation embedding $\mathbf{z}_{i}$ are updated using gradient descent: $\theta_{i}\leftarrow\theta_{i}-\eta_{\theta_{i}}\frac{\partial L_{\text{total}}}{\partial\theta_{i}}$ and $\mathbf{z}_{i}\leftarrow\mathbf{z}_{i}-\eta_{z}\frac{\partial L_{\text{total}}}{\partial\mathbf{z}_{i}}$.
+By optimizing $\mathbf{z}$, the modulation embedding learns to adjust its influence to minimize the loss, effectively enhancing inter-agent coordination and ensuring that each agent can dynamically adjust its output according to the state of the preceding agent.
+The complete training process, including the initialization of the modulation embedding, the forward pass through the agents, and the optimization of both model parameters and the modulation embedding, is detailed in Algorithm 1.
+This algorithm encapsulates our self-modulated fine-tuning approach, providing a comprehensive framework for improving the performance and coordination of our multi-agent video generation system.
+Input: Initial agents’ parameters $\{\theta_{i}\}$; initial modulation embeddings $\{\mathbf{z}_{i}=\mathcal{E}_{i}(\text{"{[Mod]}"})\}_{i=1}^{n}$; number of iterations $N$; number of samples per iteration $S$; prompt set $P$; batch size $B$; number of epochs per iteration $K$; learning rates $\eta_{\theta_{i}}$, $\eta_{z_{i}}$
+Output: Trained model parameters $\{\theta_{i}\}$
+1 for iteration $n=1$ to $N$ do
+2
+3      Construct training dataset $\mathcal{D}_{n}$ using the data-free training strategy (see Sec. 3.2 );
+4
+5      for epoch $e=1$ to $K$ do
+6             foreach batch $\{(P^{(b)},V^{(b)}_{\text{target}})\}_{b=1}^{B}$ in $\mathcal{D}_{n}$ do
+7                   Compute text embeddings $\mathbf{e}_{i}^{(b)}=\mathcal{E}_{i}(P^{(b)})$;
+8                   for $i=1$ to $n$ do
+9                         Concatenate agent-specific modulation embedding with text embedding for all examples: $\tilde{\mathbf{e}}_{i}^{(b)}=[\mathbf{e}_{i}^{(b)};\mathbf{z}_{i}]$;
+10                         Generate outputs for the batch: $O_{i}^{(b)}=\mathcal{M}_{i}(O_{i-1}^{(b)},\tilde{\mathbf{e}}_{i}^{(b)})$, where $O_{0}^{(b)}=\emptyset$;
+11
+12                  Compute the final loss for the batch: $L_{\text{final}}=\frac{1}{B}\sum_{b=1}^{B}L(O_{n}^{(b)},V^{(b)}_{\text{target}})$;
+13
+14                  for $i=1$ to $n$ do
+15                         Update agent-specific modulation embedding: $\mathbf{z}_{i}\leftarrow\mathbf{z}_{i}-\eta_{z_{i}}\frac{\partial L_{\text{final}}}{\partial\mathbf{z}_{i}}$;
+16                         Compute modulation factor: $||\mathbf{z}_{i}||_{2}=\sqrt{\sum_{k}z_{i,k}^{2}}$;
+17                         Compute modulation factor for model $i$: $\alpha_{i}=||\mathbf{z}_{i}||_{2}/n$;
+18                         Update model parameters with dynamic learning rate: $\theta_{i}\leftarrow\theta_{i}-\alpha_{i}\eta_{\theta_{i}}\frac{\partial L_{\text{final}}}{\partial\theta_{i}}$;
+19
+20
+21
+22
+23return $\{\theta_{i}\}$;
+Algorithm 1 Self-Modulated Multi-Agent Video Generation with Data-free Fine-tuning
+Multimodal LLM Selection with Human-in-the-loop Control.
+Despite the availability of numerous open-source video datasets, their quality varies significantly, making it challenging for the pretrained agents we use to effectively leverage these datasets to improve video generation performance.
+Also, manually filtering high-quality data is time-consuming and inefficient.
+To address this issue, we introduce a multimodal LLM data selection procedure with human-in-the-loop control.
+We first sample a batch of candidates from the agent system.
+Based on the generated candidates, we further leverage strong multimodal LLMs that support multi-frame and multi-video inputs to provide evaluation for the candidate set.
+When multiple multimodal LLMs agree on the best video from a set, we directly include it in the training dataset, the evaluation prompt examples detailed in Appendix A.6.
+However, when the LLMs’ evaluations differ, we introduce human reviewers for secondary screening.
+In such cases, human reviewers evaluate the generated videos and select the one with the highest quality for inclusion in the training set.
+If the reviewers determine that none of the candidates meet the quality standard, the entire set is discarded.
+This approach enhances the stability and robustness of the filtering process by combining LLMs’ automated assessments with human judgment, ensuring the model can handle complex or ambiguous cases.
+Data-free Training Strategy.
+During training, instead of using open-source datasets directly, we adopt a data-free strategy to synthesize training data dynamically.
+This method addresses challenges like inconsistent quality in open-source datasets and the lack of intermediate outputs required for training different agents.
+Given a user input prompt set $P=\{P^{(1)},P^{(2)},\dots,P^{(S)}\}$, we first utilize the LLM to generate an enhanced set of diverse prompts $P_{\text{enh}}=\{P_{\text{enh}}^{(1)},P_{\text{enh}}^{(2)},\dots,P_{\text{enh}}^{(S)}\}$.
+For each enhanced prompt $P_{\text{enh}}^{(s)}$, our initial workflow—comprising non-finetuned models with parameters $\{\theta_{i}\}_{i=1}^{M}$—is applied to generate a set of candidate videos $\mathcal{C}_{s}=\{V_{1},V_{2},V_{3},V_{4}\}$.
+Next, we use a human-in-the-loop process, in combination with multimodal LLM selection, to select the highest quality video $\hat{V}^{(s)}$ from the candidate set $\mathcal{C}_{s}$ for each prompt $P_{\text{enh}}^{(s)}$.
+The selected video $\hat{V}^{(s)}$ and corresponding prompt $P_{\text{enh}}^{(s)}$ form the training dataset for iteration $n$, denoted by $\mathcal{D}_{n}=\{(P_{\text{enh}}^{(s)},\hat{V}^{(s)})\}_{s=1}^{S}$.
+Using the constructed dataset $\mathcal{D}_{n}$, self-modulated multi-agent fine-tuning is performed to update the agent parameters $\{\theta_{i}\}_{i=1}^{M}$, where $M$ is the number of agents involved.
+The newly fine-tuned model parameters $\{\theta_{i}\}_{i=1}^{M}$ then become the starting point for the next round of data generation and fine-tuning, enabling iterative improvement over $N$ iterations.
+Thus, after $N$ iterations, the final optimized agent parameters $\{\theta_{i}\}$ are obtained.
+Figure 2: Samples for text-to-video generation of Mora.
+Our approach can generate high-resolution, temporally consistent videos from text prompts.
+The samples shown are 480p resolution over 12 seconds duration at 276 frames in total.
+Figure 3: Performance variations of Task 5 and Task 6 across different self-training iterations.
 4 Experiments
 4.1 Setup
+Tasks.
+We evaluated our proposed Mora framework on six diverse video generation tasks to comprehensively assess its capabilities.
+The six tasks are as follows: (1) Text-to-Video Generation, where videos are generated directly from textual prompts; (2) Image-to-Video Generation, which involves creating videos conditioned on both an input image and a textual description; (3) Extend Generated Videos, aiming to extend existing videos to produce longer sequences while maintaining content consistency; (4) Video-to-Video Editing, which edits existing videos based on textual instructions to produce modified versions; (5) Connect Videos, focusing on seamlessly connecting two videos to
+Figure 4: Performance variations of Task 1 to Task 4 across different self-training iterations.
+create a longer continuous sequence; and (6) Simulate Digital Worlds, where videos that simulate digital or virtual environments are generated from textual prompts.
+Data.
+For the text-to-video generation task, we utilized textual prompts inspired by those provided in the official Sora technical report (OpenAI, 2024b).
+To expand our dataset, we employed GPT-4 (OpenAI, 2023) in both few-shot and zero-shot settings to generate additional prompts.
+These prompts were then used as inputs for the text-to-video models to generate videos.
+For the other tasks, we used relevant input data appropriate for each task, such as images, videos, or textual instructions.
+For comparison with Sora, we utilized videos featured on its official website and technical report.
 Baseline.
-In the text-to-video generation, existing open-sourced works that show competitive performance are employed as the baseline models, including Videocrafter1 Chen et al.
-(2023c), Show-1 Zhang et al.
-(2023a), Pika pik , Gen-2 Gen , ModelScope Mod , LaVie-Interpolation, LaVie Wang et al.
-(2023b) and CogVideo Hong et al.
-(2022b).
+In the text-to-video generation process, we selected and compared a wide range of open-source and commercial models, including Sora (OpenAI, 2024a), VideoCrafter1 (Chen et al., 2023a), MedelScope (Mod, ), Show-1 (Zhang et al., 2023a), Pika (pik, ), LaVie and Lavie-Interpolation (Wang et al., 2023a), Gen-2 (Gen, a), and CogVideo (Hong et al., 2022a).
 In the other five tasks, we compare Mora with Sora.
-Basic Metrics.  For text-to-video generation, we employed several metrics from Vbench Huang et al. (2023) for evaluation from two aspects: video quality and video condition consistency.
-For video quality measurement, we use six metrics.
-❶ Object Consistency, computed by the DINO Caron et al.
-(2021) feature similarity across frames to assess whether object appearance remains consistent throughout the whole video; ❷ Background Consistency, calculated by CLIP Radford et al.
-(2021) feature similarity across frames; ❸ Motion Smoothness, which utilizes the motion priors in the video frame interpolation model AMT Li et al.
-(2023) to evaluate the smoothness of generated motions; ❹ Aesthetic Score, obtained by using the LAION aesthetic predictor lio on each video frame to evaluate the artistic and beauty value perceived by humans, ❺ Dynamic Degree, computed by employing RAFT Teed and Deng (2020) to estimate the degree of dynamics in synthesized videos; ❻ Imaging Quality, calculated by using MUSIQ Ke et al.
-(2021) image quality predictor trained on SPAQ Fang et al.
-(2020) dataset.
-For measuring video condition consistency, we use two metrics.
-❶ Temporal Style, which is determined by utilizing ViCLIP Wang et al.
-(2023c) to compute the similarity between video features and temporal style description features, thereby reflecting the consistency of the temporal style; ❷ Appearance Style, by calculating the feature similarity between synthesized frames and the input prompt using CLIP Radford et al.
-(2021), to gauge the consistency of appearance style.
-Self-defined Metrics.  For evaluating other tasks, we also define four metrics.
-❶ Video-Text Integration $VideoTI$, devised to enhance the quantitative evaluation of the model’s fidelity to textual instructions.
-It employs LLaVA Liu et al.
-(2024b) to transfer input image into textual descriptors $T_{i}$ and Video-Llama Zhang et al.
-(2023b) to transfer videos generated by the model into textual $T_{v}$.
-The textual representation of the image is prepended with the original instructional text, forming an augmented textual input $T_{mix}$.
-Both the newly formed text and the video-generated text will be input to BERT Devlin et al.
-(2018).
-The embeddings obtained are analyzed for semantic similarity through the computation of cosine similarity, providing a quantitative measurement of the model’s adherence to the given instructions and image.
-$VideoTI=cosine(embed_{mix},embed_{v})$ (1)
-where $embed_{mix}$ represents the embedding for $T_{mix}$ and $embed_{v}$ for $T_{v}$.
-❷ Temporal Consistency $TCON$, designed to measure the coherence between an original video and its extended version, provides a vital tool for assessing the integrity of extended video content.
-For each input-output video pair, we employ ViCLIP Wang et al.
-(2023c) video encoder to extract their feature vectors.
-We then compute cosine similarity to get the score.
-$TCON=cosine(V_{input},V_{output})$ (2)
-❸ Temporal coherence $Tmean$, by quantifying the correlation between the intermediate generated video and the input videos based on $TCON$.
-$Tmean=(TCON_{front}+TCON_{beh})/2$ (3)
-where $TCON_{front}$ measures the correlation between the intermediate video and the preceding video in the time series while $TCON_{beh}$ assesses the correlation with the subsequent video.
-The average of these scores provides an aggregate measure of temporal coherence across the video sequence.
-❹ Video Length, to evaluate the models’ efficiency in producing video content, specifically focusing on the maximum duration, measured in seconds.
+More visual comparisons and detailed analysis of the 5 tasks can be found in the Appendix A.3.
+Metrics.
+We employed a combination of standard and self-defined metrics to evaluate performance across the six tasks.
+For text-to-video generation, we employed several metrics from Vbench (Huang et al., 2024) for comprehensive evaluation from two aspects: video quality and video condition consistency.
+For video quality measurement, we use six metrics: ❶ Object Consistency, ❷ Background Consistency, ❸ Motion Smoothness, ❹ Aesthetic Score, ❺ Dynamic Degree and ❻ Imaging Quality.
+For measuring video condition consistency, we use two metrics: ❶ Temporal Style and ❷ Appearance Style.
+When evaluating other tasks, due to the lack of quantitative metrics, we design the following four metrics: ❶ Video-Text Integration $VideoTI$, ❷ Temporal Consistency $TCON$, ❸ Temporal coherence $Tmean$ and ❹ Video Length.
+For the details of these metrics, please refer to the Appendix A.2.
+4.2 Training
+In the training process, We initially used GPT-4o to generate 16 different text prompts and directly produced four videos for each prompt using an untrained model.
+During the data selection phase, we utilized two Multimodal LLMs, LLaVA-OneVision (Li et al., 2024a) and Qwen2-VL (Wang et al., 2024), to evaluate each set of videos and select the best-performing sample for each prompt.
+In the training loop, newly generated text prompts from GPT-4o were used, and after the selection process, the best videos were fed into the trained model for further optimization, and a total of 96 data were generated.
+Figures 4 and Figure 3 illustrate the impact of training on the performance of different tasks.
+It is evident that, during the early iterations, all tasks exhibit significant performance improvements, though the rate of improvement gradually slows as more iterations are completed.
+For details on the hyper-parameters settings in these training, please refer to Appendix A.9.
 Table 1: Comparative analysis of text-to-video generation performance between Mora and various other models.
 The Others category scores are derived from the Hugging Face leaderboard.
-For Our Mora, the evaluation is based on prompts generated by GPT-4, categorized into three types based on the number of moving objects in the videos: Type I (single object in motion), Type II (two to three objects in motion), and Type III (more than three objects in motion).
-Differences in the input prompt may account for the superiority of Mora inType II prompt’s scores over those of Sora in the relevant evaluations.
+For Our Mora (SVD), categorized into three types based on the number of moving objects in the videos: Type I (single object in motion), Type II (two to three objects in motion), and Type III (more than three objects in motion).
+Descriptions of Mora (SVD) and Mora (Open-Sora-Plan) are detailed in Appendix A.8.
+indicates that without Self-Modulated Multi-Agent Finetuning.
 Model Video Quality Object Consistency Background Consistency Motion Smoothness Aesthetic Quality Dynamic Dgree Imaging Quality Temporal Style Video Length(s)
@@
 Type III 0.795 0.94 0.93 0.99 0.55 0.80 0.56 0.26 12
-Mora 0.792 0.95 0.95 0.99 0.57 0.70 0.59 0.26 12
-Implementation Details.
-For text-to-video generation, we follow the style of text prompts provided in the official Sora technical report OpenAI (2024b).
-Subsequently, we employ GPT-4 OpenAI (2023b) to produce more text under a few-shot setting.
-GPT-4 is also utilized to generate the same number of texts in a zero-shot setting.
-All generated text prompts are then input into text-to-video models for generating videos.
-For comparison with Sora, we utilize videos featured on its official website and technical report.
-All experiments are conducted on two TESLA A100 GPUs, equipped with a substantial 2$\times$80GB of VRAM.
-The central processing was handled by 4×AMD EPYC 7552 48-Core Processors.
-Memory allocation was set at 320GB.
-The software environment was standardized on PyTorch version 2.0.2 and CUDA 12.2 for video generation and PyTorch version 1.10.2 and CUDA 11.6 for video evaluation.
+Mora (SVD) 0.792 0.95 0.95 0.99 0.57 0.70 0.59 0.26 12
+Mora (Open-Sora-Plan) 0.767 0.94 0.95 0.99 0.61 0.43 0.68 0.26 12
+Mora (Open-Sora-Plan) 0.800 0.98 0.97 0.99 0.66 0.50 0.70 0.31 12
 4.3 Results
@@
 The quantitative results are detailed in Table 1.
+Mora demonstrated outstanding performance in its two untrained versions, the Dynamic Degree of Mora (SVD) achieves 0.70, matching Sora and surpassing all other comparative models, which clearly highlights the effectiveness of our multi-agent collaborative architecture.
+More impressively, finetuned Mora (OSP) models outperformed all baseline methods, including Sora, achieving state-of-the-art (SoTA) performance on multiple benchmarks.
+In the following experiment content, unless otherwise specified, Mora refers to Finetuned Mora (OSP).
+Mora demonstrates exceptional performance in maintaining overall consistency in video generation.
+Specifically, it excels in the Object Consistency task, achieving a leading score of 0.98 compared to other models.
+In terms of Background Consistency, Mora scores 0.97, comparable to the state-of-the-art performance of VideoCrafter1 (Chen et al., 2023a).
+This highlights Mora’s strong capability to manage the overall coherence of generated videos.
+Mora achieved a Motion Smoothness score of 0.99, nearly matching Sora’s perfect score of 1.0, demonstrating Mora’s exceptional control over the smoothness of video sequences.
+It also outperformed in both Aesthetic Quality and Imaging Quality, with scores of 0.66 and 0.70, respectively.
+This reflects Mora’s ability to not only ensure high imaging quality but also pursue a strong aesthetic dimension in its video generation.
+Our architecture and training method not only surpass Sora in performance but also achieve this with minimal computational resource requirements, showcasing exceptional optimization efficiency.
+In Figure 2, we present examples of the generated videos.
+The visual fidelity of Mora’s text-to-video generation is compelling, manifesting high-resolution imagery with acute attention to detail as articulated in the accompanying textual descriptions.
+Notably, the images exude a temporal consistency that speaks to Mora’s nuanced understanding of narrative progression, an essential quality in video synthesis from textual prompts.
+Image-to-Video Generation.
+Quantitative comparisons between Sora and Mora across different tasks are presented in Table 2.
+As shown in the table, Mora demonstrates comparable performance as Sora in the Motion Smoothness metric, achieving 0.98.
+In the VideoTI metric, both Mora and Sora are tied at 0.90.
+For the remaining metrics, Mora surpasses all other comparative methods, particularly achieving a perfect score of 1.0 in Dynamic Degree, demonstrating its exceptional capability in enhancing the sense of motion in videos.
+Additionally, Mora leads significantly in image quality with a score of 0.67, clearly indicating its superior performance in image rendering within video sequences.
+Also further demonstrates the usability and strong performance of Mora in the image-to-video task.
+Extend Generated Video.
+The quantitative results from Table 2 reveal that Mora demonstrates similar performance as Sora in TCON and Temporal Style, with scores of 0.98 and 0.23 compared to Mora’s 0.99 and 0.24.
+This indicates that Mora is nearly on par with Sora in maintaining stylistic continuity as well as sequence consistency and fidelity.
+In terms of Imaging Quality, Mora surpasses all other methods with scores of 0.45, demonstrating its excellent imaging capabilities.
+Despite Sora’s very narrow lead, Mora effectively extends videos while maintaining high imaging quality, adhering to temporal style, and ensuring strong stylistic continuity and sequence consistency, highlighting its effectiveness in the video extension domain.
+Video-to-Video Editing  Table 2 shows that Mora achieves respective scores of 0.52 and 0.24 in Imaging Quality and Temporal Style, respectively, matching Sora in both aspects.
+This indicates that Mora not only consistently achieves high-quality imaging in video editing but also effectively maintains stylistic consistency throughout.
+This also further highlights Mora’s usability and strong adaptability in video-to-video editing tasks.
+Connect Videos.
+Quantitative analysis in Table 2 shows that Sora outperforms Mora in both Imaging Quality and Tmean.
+Sora scores 0.52 in Imaging Quality and 0.64 in Temporal Coherence, while Mora’s best scores are 0.43 and 0.45, respectively.
+This demonstrates Sora’s superior fidelity in visual representation and consistency in visual narrative.
+These results highlight Sora’s advantage in creating high-quality, temporally coherent video sequences, while also indicating areas where Mora could be further improved.
+Simulate Digital Worlds.
+Table 2 shows that Sora leads in digital world simulation with a higher Imaging Quality score of 0.62, indicating better visual realism and fidelity compared to Mora’s score of 0.52.
+However, both models achieve identical scores in Appearance Style at 0.23, indicating they equally adhere to the stylistic parameters of the digital worlds.
+This suggests that while there is a difference in the imaging quality, the stylistic translation of textual descriptions into visual aesthetics is accomplished with equivalent proficiency by both models.
+More results, examples, ablation studies, and case studies are provided in Appendix B.
+Table 2: Summary of model evaluations across various tasks based on the Sora technical report (OpenAI, 2024b).
+Model Image-to-Video Generation Extend Videos
+VideoTI Motion Smoothness Dynamic Degree Imaging Quality TCON Imaging Quality Temporal Style
+Sora 0.90 0.99 0.75 0.63 0.99 0.43 0.24
+SVD 0.88 0.97 0.75 0.58 0.94 0.37 0.22
+SVD-1.1 0.88 0.97 0.75 0.59 0.94 0.37 0.22
+Open-Sora-Plan 0.89 0.98 1.00 0.65 0.96 0.45 0.22
+Mora (SVD) 0.88 0.97 0.75 0.60 0.94 0.39 0.22
+Mora (Open-Sora-Plan) 0.88 0.97 1.00 0.66 0.96 0.45 0.22
+Mora (Open-Sora-Plan) 0.90 0.98 1.00 0.67 0.98 0.45 0.23
+Model Video-to-Video Editing Connect Videos Simulate Digital Worlds
+Imaging Quality Temporal Style Imaging Quality Tmean Imaging Quality Appearance Style
+Sora 0.52 0.24 0.52 0.64 0.62 0.23
+Pika 0.35 0.20 - - 0.44 0.10
+Runway-tool - - 0.33 0.22 - -
+Mora (SVD) 0.38 0.23 0.42 0.45 0.52 0.23
+Mora (Open-Sora-Plan) 0.34 0.20 0.40 0.39 0.51 0.20
+Mora (Open-Sora-Plan) 0.52 0.24 0.43 0.44 0.56 0.23
+5 Conclusion
+We introduce Mora, a pioneering generalist framework that synergies Standard Operating Procedures (SOPs) for video generation, tackling a broad spectrum of video-related tasks.
+Mora significantly advances the generation of videos from textual prompts, establishing new standards for adaptability, efficiency, and output quality in the domain.
+Our comprehensive evaluation indicates that Mora not only matches but also surpasses the capabilities of existing leading models in several respects.
+Nonetheless, it still faces a significant performance gap compared to OpenAI’s Sora, whose closed-source nature presents substantial challenges for replication and further innovation in both academic and professional settings.
+Looking forward, there are multiple promising avenues for further research.
+One direction includes enhancing the natural language understanding capabilities of the agents to support more nuanced and context-aware video productions.
+Additionally, the issues of accessibility and high computational demands continue to impede broader adoption and innovation.
+Concurrently, fostering more open and collaborative research environments could spur advancements, allowing the community to build upon the foundational achievements of the Mora framework and other leading efforts.
+Appendix A Implementation Details
+A.1 Hardware Details
+All experiments are conducted on eight TESLA H100 GPUs, equipped with a substantial 8$\times$80GB of VRAM.
+The central processing was handled by 4×AMD EPYC 7552 48-Core Processors.
+Memory allocation was set at 320GB.
+The software environment was standardized on PyTorch version 2.0.2 and CUDA 12.2 for video generation and one TESLA A40, PyTorch version 1.10.2 and CUDA 11.6 for video evaluation.
+A.2 Metrics
+Basic Metrics.
+❶ Object Consistency, computed by the DINO (Caron et al., 2021) feature similarity across frames, assesses whether object appearance remains consistent throughout the entire video; ❷ Background Consistency, calculated by CLIP (Radford et al., 2021) feature similarity across frames; ❸ Motion Smoothness, utilizing motion priors in the video frame interpolation model AMT (Li et al., 2023) to evaluate the smoothness of generated motions; ❹ Aesthetic Score, obtained by using the LAION aesthetic predictor (lio, ) on each video frame to evaluate the artistic and beauty value perceived by humans, ❺ Dynamic Degree, computed by employing RAFT (Teed & Deng, 2020) to estimate the degree of dynamics in synthesized videos; ❻ Imaging Quality, calculated by MUSIQ (Ke et al., 2021) image quality predictor trained on SPAQ (Fang et al., 2020) dataset.
+❶ Temporal Style, which is determined by utilizing ViCLIP (Wang et al., 2023b) to compute the similarity between video features and temporal style description features, thereby reflecting the consistency of the temporal style; ❷ Appearance Style, by calculating the feature similarity between synthesized frames and the input prompt using CLIP (Radford et al., 2021), to gauge the consistency of appearance style.
+Self-defined Metrics.
+❶ Video-Text Integration ($VideoTI$) employs LLaVA (Liu et al., 2024) to transfer input image into textual descriptors $T_{i}$ and Video-Llama (Zhang et al., 2023b) to transfer videos generated by the model into textual $T_{v}$.
+The textual representation of the image is prepended with the original instructional text, forming an augmented textual input $T_{mix}$.
+Both the newly formed text and the video-generated text will be input to BERT (Devlin et al., 2018).
+The embeddings obtained are analyzed for semantic similarity through the computation of cosine similarity, providing a quantitative measurement of the model’s adherence to the given instructions and image.
+$VideoTI=cosine(embed_{mix},embed_{v})$ (1)
+where $embed_{mix}$ represents the embedding for $T_{mix}$ and $embed_{v}$ for $T_{v}$.
+❷ Temporal Consistency ($TCON$) assesses the integrity of extended video content.
+For each input-output video pair, we employ ViCLIP (Wang et al., 2023b) video encoder to extract their feature vectors.
+We then compute cosine similarity to get the score.
+$TCON=cosine(V_{input},V_{output})$ (2)
+❸ Temporal coherence $Tmean$ evaluates the temporal coherence by calculating the average correlation between intermediate generated videos and their neighboring input videos.
+Specifically, $TCON_{front}$ measures the correlation between the intermediate video and the preceding video in the time series, while $TCON_{beh}$ assesses the correlation with the subsequent video.
+The average of these scores provides an aggregate measure of temporal coherence across the video sequence.
+$Tmean=(TCON_{front}+TCON_{beh})/2$ (3)
+A.3 Details of SOPs
+Text-to-Video Generation.
+This task harnesses a detailed textual prompt from the user as the foundation for video creation.
+The prompt must meticulously detail the envisioned scene.
+Utilizing this prompt, the Text-to-Image agent utilizes the text, distilling themes and visual details to craft an initial frame.
+Building upon this foundation, the Image-to-Video component methodically generates a sequence of images.
+This sequence dynamically evolves to embody the prompt’s described actions or scenes, and each video is derived from the last frame from the previous video, thereby achieving a seamless transition throughout the video.
+Image-to-Video Generation.
+This task mirrors the operational pipeline of Task 1, with a key distinction.
+Unlike Task 1 with only texts as inputs, Task 2 integrates both a textual prompt and an initial image into the Text-to-Image agent’s input.
+This dual-input approach enriches the content generation process, enabling a more nuanced interpretation of the user’s vision.
+Extend Generated Videos.
+This task focuses on extending the narrative of an existing video sequence.
+By taking the last frame of an input video as the starting point, the video generation agent crafts a series of new, coherent frames that continue the story.
+This approach allows for the seamless expansion of video content, creating longer narratives that maintain the consistency and flow of the original sequence without disrupting its integrity.
+Video-to-Video Editing.
+This task introduces a sophisticated editing capability, leveraging both the Image-to-Image and Image-to-Video agents.
+The process begins with the Image-to-Image agent, which takes the first frame of an input video and applies edits based on the user’s prompt, achieving the desired modifications.
+This edited frame then serves as the initial image for the Image-to-Video agent, which generates a new video sequence that reflects the requested obvious or subtle changes, offering a powerful tool for dynamic video editing.
+Connect Videos.
+The Image-to-Video agent leverages the final frame of the first input video and the initial frame of the second input video to create a seamless transition, producing a new video that smoothly connects the two original videos.
+Simulating Digital Worlds.
+This task specializes in the whole style changing for video sequences set in digitally styled worlds.
+By appending the phrase "In digital world style" to the edit prompt, the user instructs the Image-to-Video agent to craft a sequence that embodies the aesthetics and dynamics of a digital realm or utilize the Image-to-Image agent to transfer the real image to digital style.
+This task pushes the boundaries of video generation, enabling the creation of immersive digital environments that offer a unique visual experience.
+Figure 5: An example of image generation process in Mora.
+Left: Agent uses the structured message to communicate, Right: After the prompt or image is generated, a human user can check the quality of the generated content.
+A.4 Iterative Programming with Human in the Loop
+Human oversight and iterative refinement are essential in content generation tasks, improving the quality and precision of the final outputs.
+Integrating human collaboration within video generation frameworks is pivotal for ensuring that the resulting videos conform to the expected standards.
+As illustrated in Figure 1 and 5, Mora is engineered to generate videos based on input prompts while executing SOPs under continuous human supervision.
+This supervisory role enables users to rigorously monitor the process and verify that the outputs align with their expectations.
+Following each stage, users have the discretion to either advance to the subsequent phase or request a repetition of the previous one, with a maximum of three iterations allowed per stage.
+This iterative mechanism persists until the output is deemed satisfactory, ensuring high fidelity to user requirements.
+A.5 Function of Agents
+Prompt Selection and Generation Agent.
+Prior to the commencement of the initial image generation, textual prompts undergo a rigorous processing and optimization phase.
+This critical agent can employ large language models like GPT-4, GPT-4o, Llama, Llama3 (Achiam et al., 2023; GPT, ; Touvron et al., 2023).
+It is designed to meticulously analyze the text, extracting pivotal information and actions delineated within, thereby significantly enhancing the relevance and quality of the resultant images.
+This step ensures that the textual descriptions are thoroughly prepared for an efficient and effective translation into visual representations.
+Text-to-Image Generation Agent.
+The text-to-image model (Rombach et al., 2022b; Podell et al., 2023) stands at the forefront of translating these enriched textual descriptions into high-quality initial images.
+Its core functionality revolves around a deep understanding and visualization of complex textual inputs, enabling it to craft detailed and accurate visual counterparts to the provided textual descriptions.
+Image-to-Image Generation Agent.
+This agent (Brooks et al., 2023) works to modify a given source image in response to specific textual instructions.
+The core of its functionality lies in its ability to interpret detailed textual prompts with high accuracy and subsequently apply these insights to manipulate the source image accordingly.
+This involves a detailed recognition of the text’s intent, translating these instructions into visual modifications that can range from subtle alterations to transformative changes.
+The agent leverages a pre-trained model to bridge the gap between textual description and visual representation, enabling seamless integration of new elements, adjustment of visual styles, or alteration of compositional aspects within the image.
+Image-to-Video Generation Agent.
+Following the creation of the initial image, the Video Generation Model (Blattmann et al., 2023) is responsible for transitioning the static frame into a vibrant video sequence.
+This component delves into the analysis of both the content and style of the initial image, serving as the foundation for generating subsequent frames.
+These frames are meticulously crafted to ensure a seamless narrative flow, resulting in a coherent video that upholds temporal stability and visual consistency throughout.
+This process highlights the model’s capability to not only understand and replicate the initial image but also to anticipate and execute logical progressions in the scene.
+Video Connection Agent.
+Utilizing the Video-to-Video Agent, we create seamless transition videos based on two input videos provided by users.
+This advanced agent selectively leverages key frames from each input video to ensure a smooth and visually consistent transition between them.
+It is designed with the capability to accurately identify the common elements and styles across the two videos, thus ensuring a coherent and visually appealing output.
+This method not only improves the seamless flow between different video segments but also retains the distinct styles of each segment.
+A.6 MLLMs to Judge Video Quality
+Prompt No. Evaluation Criteria
+1 Evaluate the visual clarity and resolution, ranking videos based on image sharpness, smoothness of transitions, and noise levels.
+2 Assess object consistency and scene stability across frames, ranking videos on object motion and interactions.
+3 Examine the temporal coherence, identifying the best frame-to-frame continuity.
+4 Evaluate the narrative coherence or logical progression, ranking based on storytelling consistency.
+5 Assess color grading and lighting consistency, determining the best video based on smooth lighting transitions and uniform color.
+6 Evaluate the realism of objects, background textures, and scene complexity, ranking videos from most realistic to least.
+7 Analyze content relevance to the task, ranking videos based on theme alignment and task appropriateness.
+8 Compare the aesthetic quality, focusing on artistic composition, balance, and overall visual appeal.
+9 Evaluate noise and artifact levels, identifying the video with the cleanest and smoothest output.
+10 Examine frame rate consistency and smoothness of motion, ranking videos based on natural motion without lag or stuttering.
+Table 3: Evaluation Prompts for MLLMs to Judge Video Quality
+We employ two multimodal large language models (MLLMs) to evaluate four videos, using random prompts selected from Table 3.
+The MLLMs process the inputs and generate rankings based on the specified evaluation criteria.
+From the output, we focus on analyzing the video ranked as top-1 by each MLLM.
+If both MLLMs agree on the top-1 video, it is directly included in the training dataset.
+In cases where the MLLMs disagree, human reviewers intervene to further assess the videos and finalize the selection.
+This hybrid approach ensures the inclusion of high-quality, relevant videos, while balancing automation with human judgment.
+By integrating MLLMs with human oversight, we maintain robustness and precision in the data selection process, which ultimately enhances the video generation model’s performance.
+A.7 Details of Training Pipeline
+Figure 6: Illustration of the process of constructing training dataset and the design of our self-modulated fine-tuning.
+The training pipeline is illustrated in two key phases, as shown in Figure 6, constructing the training dataset and self-modulated fine-tuning.
+Constructing the Training Dataset.
+The process starts with user input prompts that are enhanced by a multimodal large language model (MLLM).
+These enhanced prompts are evaluated using a human-in-the-loop strategy, which ensures high-quality video candidates.
+The selected videos, paired with their prompts, form the final training data for fine-tuning the model.
+Self-Modulated Fine-Tuning.
+During fine-tuning, enhanced prompts are processed by multiple agents.
+Each agent generates text embeddings, which are combined with a modulation embedding to dynamically adjust the contribution of each agent.
+The agents work together in a coordinated manner to generate video outputs.
+The total loss is calculated based on the output video and the target, and gradient descent is used to update both the model parameters and the modulation embeddings, and then reconstruct the training datasets.
+This ensures that the agents adapt their outputs based on the preceding agent’s state, resulting in improved video generation.
+A.8 Details of Agents
+Prompt Selection and Generation.
+Currently, GPT-4o  (GPT, ) stands as the most advanced generative model available.
+By harnessing the capabilities of GPT-4o, we are able to generate and meticulously select high-quality prompts.
+These prompts are detailed and rich in information, facilitating the Text-to-Image generation process by providing the agent with comprehensive guidance.
+Text-to-Image Generation.  We utilize the pre-trained large text-to-image model to generate a high-quality and representative first image.
+For Mora (SVD), the initial implementation utilizes Stable Diffusion XL (SDXL) (Podell et al., 2023).
+It introduces a significant evolution in the architecture and methodology of latent diffusion models (Rombach et al., 2022b; Meng et al., 2022) for text-to-image synthesis, setting a new benchmark in the field.
+At the core of its architecture is an enlarged UNet backbone (Ronneberger et al., 2015) that is three times larger than those used in previous versions of Stable Diffusion 2 (Rombach et al., 2022b).
+This expansion is principally achieved through an increased number of attention blocks and a broader cross-attention context, facilitated by integrating a dual text encoder system.
+The first encoder is based on OpenCLIP (Ilharco et al., 2021) ViT-bigG (Cherti et al., 2023; Radford et al., 2021; Schuhmann et al., 2022), while the second utilizes CLIP ViT-L, allowing for a richer, more nuanced interpretation of textual inputs by concatenating the outputs of these encoders.
+This architectural innovation is complemented by the introduction of several novel conditioning schemes that do not require external supervision, enhancing the model’s flexibility and capability to generate images across multiple aspect ratios.
+Moreover, SDXL features a refinement model that employs a post-hoc image-to-image transformation to elevate the visual quality of the generated images.
+This refinement process utilizes a noising-denoising technique, further polishing the output images without compromising the efficiency or speed of the generation process.
+For Mora (Open-Sora-Plan), the Stable Diffusion 3 (SD3) (Esser et al., 2024) text-to-image model employed to generate high-quality images from textual prompts.
+SD3 utilizes a rectified flow transformer architecture, offering significant advancements over traditional diffusion models by connecting data and noise in a linear path rather than the curved trajectories commonly used in previous approaches.
+This architectural choice facilitates a more efficient and accurate sampling process, enhancing the generation of high-resolution images across diverse styles.
+Additionally, SD3 introduces a novel noise re-weighting technique that biases sampling toward perceptually relevant scales, thereby improving the clarity and aesthetic quality of the generated images.
+The model also supports multiple aspect ratios and incorporates a refinement process that further enhances generated images through a noising-denoising technique, ensuring high visual fidelity while maintaining computational efficiency.
+Image-to-Image Generation.
+Our initial Mora (SVD) framework utilizes InstructPix2Pix as Image-to-Image generation agent.
+InstructPix2Pix  (Brooks et al., 2023) are intricately designed to enable effective image editing from natural language instructions.
+At its core, the system integrates the expansive knowledge of two pre-trained models: GPT-3  (Brown et al., 2020) for generating editing instructions and edited captions from textual descriptions, and Stable Diffusion  (Rombach et al., 2022a) for transforming these text-based inputs into visual outputs.
+This ingenious approach begins with fine-tuning GPT-3 on a curated dataset of image captions and corresponding edit instructions, resulting in a model that can creatively suggest plausible edits and generate modified captions.
+Following this, the Stable Diffusion model, augmented with the Prompt-to-Prompt technique, generates pairs of images (before and after the edit) based on the captions produced by GPT-3.
+The conditional diffusion model at the heart of InstructPix2Pix is then trained on this generated dataset.
+InstructPix2Pix directly utilizes the text instructions and input image to perform the edit in a single forward pass.
+This efficiency is further enhanced by employing classifier-free guidance for both the image and instruction conditionings, allowing the model to balance fidelity to the original image with adherence to the editing instructions.
+For Mora (Open-Sora-Plan), the image-to-image generation agent, based on SD3 (Esser et al., 2024), excels in applying detailed text-guided transformations to images with precision and flexibility.
+SD3’s rectified flow transformer architecture ensures efficient and accurate image modifications, while the noise re-weighting process enhances the visual quality of the output, ensuring that edits are seamless and coherent.
+Image-to-Video Generation.  In the Text-to-Video generation agent, video generation agents play an important role in ensuring video quality and consistency.
+Our initial Mora (SVD) implementation utilizes the state-of-the-art video generation model Stable Video Diffusion to generate video.
+The Stable Video Diffusion (SVD)  (Blattmann et al., 2023) architecture introduces a cutting-edge approach to generating high-resolution videos by leveraging the strengths of LDMs Stable Diffusion v2.1  (Rombach et al., 2022a), originally developed for image synthesis, and extending their capabilities to handle the temporal complexities inherent in video content.
+At its core, the SVD model follows a three-stage training regime that begins with text-to-image pertaining, where the model learns robust visual representations from a diverse set of images.
+This foundation allows the model to understand and generate complex visual patterns and textures.
+In the second stage, video pretraining, the model is exposed to large amounts of video data, enabling it to learn temporal dynamics and motion patterns by incorporating temporal convolution and attention layers alongside its spatial counterparts.
+This training is conducted on a systematically curated dataset, ensuring the model learns from high-quality and relevant video content.
+The final stage, high-quality video finetuning, focuses on refining the model’s ability to generate videos with increased resolution and fidelity, using a smaller but higher-quality dataset.
+This hierarchical training strategy, complemented by a novel data curation process, allows SVD to excel in producing state-of-the-art text-to-video and image-to-video synthesis with remarkable detail, realism, and coherence over time.
+The Mora (Open-Sora-Plan) video generation agent leverages a 3D full attention mechanism, which processes spatial and temporal features simultaneously, providing a unified understanding of motion and appearance across frames.
+This mechanism ensures smoother transitions and higher coherence in movements across consecutive frames, improving the temporal consistency of the generated video.
+In the second phase, temporal convolution layers are integrated to further refine the model’s capacity to comprehend and generate realistic motion dynamics.
+By capturing temporal patterns over time, the model produces videos with continuous and natural movements, thereby enhancing the realism of complex visual sequences.
+The third phase involves utilizing the CausalVideoVAE  (Chen et al., 2024b), which enhances the visual representation and detail of each frame while minimizing the occurrence of artifacts commonly associated with video synthesis.
+The CausalVideoVAE maintains high-quality visual outputs throughout the entire video sequence by refining and polishing the generated content through multiple iterations.
+This comprehensive training regime ensures that Mora (Open Sora Plan) excels in both image-to-video and text-to-video tasks, delivering exceptional levels of detail, realism, and temporal coherence.
+These architectural innovations enable superior handling of complex motion patterns and high-resolution video outputs, resulting in visually consistent and temporally stable videos.
+Connect Videos.
+For the video connection task, we utilize SEINE (Chen et al., 2023c) to connect videos.
+SEINE is constructed upon a pre-trained diffusion-based T2V model, LaVie (Wang et al., 2023a) agent.
+SEINE centered around a random-mask video diffusion model that generates transitions based on textual descriptions.
+By integrating images of different scenes with text-based control, SEINE produces transition videos that maintain coherence and visual quality.
+Additionally, the model can be extended for tasks such as image-to-video animation and autoregressive video prediction.
+In Mora (Open-Sora-Plan), the video connection task utilizes a specialized architecture to ensure seamless and visually coherent transitions between videos.
+This is achieved by combining diffusion-based models with temporal convolution techniques, facilitating smooth transitions in style, content, and motion dynamics.
+The video connection agent employs CausalVideoVAE, optimized for temporal and spatial consistency, enhancing the connection of two video segments by identifying and preserving common visual elements across frames.
+A 3D full attention architecture is central to this process, enabling the model to simultaneously capture spatial and temporal features, thereby maintaining coherence in object motion and background continuity.
+Additionally, the architecture incorporates random-mask video diffusion, which maintains high-resolution transitions by infilling missing information based on text-based control inputs or video context.
+This approach ensures that the connected videos preserve visual quality and exhibit coherent motion patterns, resulting in high-quality, temporally stable transitions.
+A.9 Details Settings in Training
+In our training setup, we use the AdamW optimizer, which is known for handling weight decay effectively, with an initial learning rate of 1e-5.
+The learning rate remains constant throughout training, controlled by the constant scheduler, and no warmup steps are used (lr_warmup_steps=0).
+To handle memory constraints, we accumulate gradients over 1 step, which allows us to use a batch size of 4 while maintaining stable optimization.
+We enable gradient checkpointing to further save memory by reducing the number of intermediate activations stored during backpropagation, although this comes at the cost of slower backward passes.
+Mixed precision training with bf16 is employed to enhance computational speed and lower memory usage, which is crucial when training with a batch size of 4.
+The model is trained for a maximum of 12 steps, with checkpoints being saved every 4 steps to allow for model recovery or evaluation during training.
+We use an SNR Gamma value of 5.0 to balance the noise scale, which is especially useful for diffusion-based models, and apply Exponential Moving Average (EMA) from step 0 with a decay rate of 0.999 to ensure model stability throughout training.
+These settings provide a balance between computational efficiency and model performance, ensuring that we can handle large video data with high memory demands while optimizing effectively across training iterations.
+Since the OpenSora and SD3 environments are mutually exclusive (due to certain package incompatibilities), we use inter-process communication (IPC) to perform joint training.
+Listing 1: SD3 Process (Text-to-Image) Pseudocode
+# Initialize ZeroMQ context and create a REQ socket
+Initialize context
+Create REQ socket
+Connect to Open-Sora-Plan environment
+# Main loop
+while training is not complete:
+# Generate image from Stable-Diffusion3 with prompt
+image = Stable_Diffusion3(prompt)
+# Send image data to Open-Sora-Plan
+socket.send(image)
+# Receive video loss from Open-Sora-Plan
+video_loss = socket.recv()
+# Backpropagation
+optimizer.zero_grad()
+video_loss.backward()
+optimizer.step()
+Listing 2: Open-Sora-Plan Process (Image-to-Video) Pseudocode
+# Initialize ZeroMQ context and create a REP socket
+Initialize context
+Create REP socket
+Bind to specific port
+# Main loop
+while True:
+# Receive image data from SD3
+image_data = socket.recv()
+# Process the image data
+image = process_image(image_data)
+# Generate video from image using Open-Sora-Plan
+generated_video = Open_Sora_Plan(image, prompts)
+# Compute video loss
+video_loss = compute_loss(generated_video, ground_truth_video)
+# Send video loss back to SD3
+socket.send(video_loss)
+Appendix B More Results and Examples
+B.1 More Results Details
+Figure 7: Comparative analysis of text-to-video generation performance between Mora and various other models.
+Text-to-Video Generation.
+The quantitative results are detailed in Table 1 and Figure 7.
 Mora showcases commendable performance across all metrics, making it highly comparable to the top-performing model, Sora, and surpassing the capabilities of other competitors.
@@
 Although this performance signifies a commendable proficiency, it also highlights a considerable gap between our model and Sora, the leader in this category with a score of 0.35.
+The results in Table 4 compare various models on video generation performance using the Vbench dataset.
+Mora (Open-Sora-Plan) achieves the highest overall video quality (0.848) and excels in both object and background consistency (0.98), while also leading in aesthetic quality (0.70) and imaging quality (0.72).
+Emu3 and Gen-3-Alpha perform similarly in video quality (0.841), with Emu3 showing superior motion smoothness (0.99) and dynamic degree (0.79).
+Despite a slightly lower video quality, LaVie-2 outperforms other models in temporal style and maintains strong consistency metrics across the board.
 In Figure 2, the visual fidelity of Mora’s text-to-video generation is compelling, manifesting high-resolution imagery with acute attention to detail as articulated in the accompanying textual descriptions.
@@
 Notably, the images exude a temporal consistency that speaks to Mora’s nuanced understanding of narrative progression, an essential quality in video synthesis from textual prompts.
+Figure 8: Demo of Mora on various tasks.
+Mora can deal with various tasks, including video editing, video extension, and stimulate digital world. We provide demo of Mora in Figure 8.
 Text-conditional Image-to-Video Generation.
-Analyzing Table 3, we discern a notable demonstration of Mora’s capabilities in text-conditional image-to-video generation, closely trailing behind Sora.
-Sora leads with a VideoTI score of 0.90 and Motion Smoothness of 0.99, underscoring its refined alignment of video with the provided text and the fluidity of motion within its generated videos.
-Mora’s VideoTI score of 0.88 and Motion Smoothness of 0.97, while marginally lower, reflect its robust potential in the accurate interpretation of textual prompts and the generation of smooth motion sequences.
-Both models exhibit a matched score in Dynamic Degree at 0.75, indicating a comparable proficiency in animating videos with a sense of activity and movement.
-The Imaging Quality scores show Sora with a slight lead at 0.63 against Mora’s 0.60, suggesting Sora’s subtle superiority in image rendering within video sequences.
-Table 3: Text-condtional image-to-video generation.
-This task evaluates the model’s ability to convert static images and accompanying textual instructions into coherent video sequences.
-We used 4 pairs of image and text inputs sourced from the official Sora technical report OpenAI (2024b).
-Model VideoTI Motion Smoothness Dynamic Dgree Imaging Quality
-Sora 0.90 0.99 0.75 0.63
-Mora 0.88 0.97 0.75 0.60
+In Figure 14, a qualitative comparison between the video outputs from Sora and Mora reveals that both models adeptly incorporate elements from the input prompt and image.
+The monster illustration and the cloud spelling "SORA" are well-preserved and dynamically translated into video by both models.
+Despite quantitative differences, the qualitative results of Mora nearly rival those of Sora, with both models are able to animate the static imagery and narrative elements of the text descriptions into coherent video.
+This qualitative observation attests to Mora’s capacity to generate videos that closely parallel Sora’s output, achieving a high level of performance in rendering text-conditional imagery into video format while maintaining the thematic and aesthetic essence of the original inputs.
 Figure 9: Samples for text-conditional image-to-video generation of Mora and Sora.
@@
 The second image’s prompt is: An image of a realistic cloud that spells "SORA".
-In Figure 3, a qualitative comparison between the video outputs from Sora and Mora reveals that both models adeptly incorporate elements from the input prompt and image.
-The monster illustration and the cloud spelling "SORA" are well-preserved and dynamically translated into video by both models.
-Despite quantitative differences, the qualitative results of Mora nearly rival those of Sora, with both models are able to animate the static imagery and narrative elements of the text descriptions into coherent video.
-This qualitative observation attests to Mora’s capacity to generate videos that closely parallel Sora’s output, achieving a high level of performance in rendering text-conditional imagery into video format while maintaining the thematic and aesthetic essence of the original inputs.
+Table 4: Comparative analysis of text-to-video generation performance for models with Vbench data.
+Model Video Quality Object Consistency Background Consistency Motion Smoothness Aesthetic Quality Dynamic Degree Imaging Quality Temporal Style Video Length(s)
+Emu3 0.841 0.95 0.98 0.99 0.60 0.79 0.63 0.24 5
+Open-Sora-Plan_V1.2 0.814 0.97 0.98 0.99 0.59 0.42 0.57 0.25 8
+Gen-3-Alpha 0.841 0.97 0.97 0.99 0.63 0.60 0.67 0.25 10
+LaVie-2 0.832 0.98 0.98 0.98 0.68 0.31 0.70 0.25 3
+Mora (Open-Sora-Plan) 0.848 0.98 0.98 0.99 0.70 0.72 0.72 0.25 12
 Extend Generated Videos.
-The quantitative results from Table 4 reveal that while Sora holds a slight edge over Mora in TCON and Imaging Quality, indicating a higher consistency and fidelity in extending video sequences, Mora’s performance remains close, with a marginal difference of 0.05 in TCON and 0.04 in Imaging Quality.
-The closeness in Temporal Style scores, 0.24 for Sora and 0.22 for Mora, further signifies that Mora nearly matches Sora in maintaining stylistic continuity over time.
-Despite Sora’s lead, Mora’s capabilities, particularly in following the temporal style and extending existing videos without significant quality loss, demonstrate its effectiveness in the video extension domain.
-Table 4: Extend generated videos.
-This experiment compares the continuity and quality of the extended video with the same input video given to Sora and Mora respectively.
-The input video is taken from the full video of the offical Sora technical report OpenAI (2024b), and we take 2 clips as the input video for this task.
-Model TCON Imaging Quality Temporal Style
-Sora 0.99 0.43 0.24
-Mora 0.94 0.39 0.22
-Figure 4: Samples for Extend generated video of Mora and Sora.
 From a qualitative standpoint, Figure 10 illustrates the competencies of Mora in extending video sequences.
@@
 This visual assessment underscores Mora’s proficiency in generating extended video content that closely mirrors the original, maintaining the narrative context and visual integrity, thus providing near parity with Sora’s performance.
-Table 5: Video-to-video editing.
-This task examines the model’s ability to edit and transform existing video content in accordance with textual instructions while maintaining visual and stylistic coherence.
-Data includes 1 input video and 12 instructions from official Sora technical report.
-Model Imaging Quality Temporal Style
-Sora 0.52 0.24
-Mora 0.38 0.23
+Figure 10: Samples for Extend generated video of Mora and Sora.
 Video-to-Video Editing.
-Table 5 exhibits a comparative analysis of video-to-video editing capabilities between Sora and Mora.
-Sora secures a higher score in Imaging Quality at 0.52, which suggests a superior capability in preserving the visual details and overall image fidelity during the video editing process.
-Mora demonstrates capability in video-to-video editing with an Imaging Quality score of 0.38.
-Although this score reveals a noticeable discrepancy when compare to Sora, it offers valuable insights on pinpointing areas for targeted enhancements in future iterations of the Mora.
-In the aspect of Temporal Style, both models exhibit proximal performance, with Sora marginally leading at 0.24 compared to Mora’s 0.23.
-This near-parity underscores the capacity of Mora to closely emulate the stylistic temporal consistency achieved by Sora, ensuring that the edited videos retain a coherent style over time, a crucial aspect of seamless video-to-video editing.
-Figure 5: Samples for Video-to-video editing
-Table 6: Connect Videos.
-This task assess the capability of models to seamlessly integrate distinct video clips into a cohesive sequence.
-$*{}^{*}$ora data we utilize is 5 distinct video-video pairs from Sora technical report website and Mora data we utilize 5 video-video pairs of similar styles.
-Model Imaging Quality Tmean
-$*{}^{*}$ 0.52 0.64
-Mora $*{}^{*}$ 0.42 0.45
-Figure 6: Samples for Connect Videos
 Upon qualitative evaluation, Figure 11 presents samples from video-to-video editing tasks, wherein both Sora and Mora were instructed to modify the setting to the 1920s style while maintaining the car’s red color.
@@
 This capability, although not as refined as Sora’s, demonstrates Mora’s potential for significant video editing tasks.
+Figure 11: Samples for Video-to-video editing
+Figure 12: Samples for Video Connetion
 Connect Videos.
-Quantitative analysis based on the results shown in Table 6 suggests that the model Sora outperforms Mora in terms of Imaging Quality and Tmean.
-Sora achieves a score of 0.52 in Imaging Quality, indicating its higher fidelity in visual representation compared to Mora’s 0.42.
-Furthermore, Sora’s superiority is evident in the Temporal coherence aspect with a score of 0.64, which implies that Sora maintains a more consistent visual narrative over time than Mora, which scores 0.45.
-This quantitative assessment not only solidifies Sora’s position as a superior model in creating high-quality, temporally coherent video sequences but also delineates a trajectory for future enhancements in video connectivity for the Mora framework.
 Qualitative analysis based on Figure 12 suggest that, in comparison to Sora’s proficiency in synthesizing intermediate video segments that successfully incorporate background elements from preceding footage and distinct objects from subsequent frames within a single frame, the Mora model demonstrates a blurred background in the intermediate videos, which results in indistinguishable object recognition.
@@
 This would contribute to refining the video connecting process and improving the integration quality of Mora’s model outputs.
-Table 7: Simulate digital worlds.
-This task assesses model’s effectiveness in creating videos that emulate digital or virtual environments, with a focus on preserving the visual details and the distinct appearance style of such worlds.
-By referring to the video generated by Sora from the official technical report OpenAI (2024b), we used GPT-4 OpenAI (2023b) to build the input prompt for Mora, which contains the "minecraft scene" and the "player character".
-Model Imaging Quality Appearance style
-Sora 0.62 0.23
-Mora 0.52 0.23
 Figure 13: Samples for Simulate digital worlds
 Simulate Digital Worlds.
-In the assessment of digital world simulation capabilities, Table 7 presents a comparative metric analysis between Sora and Mora.
-Sora exhibits a lead in Imaging Quality with a score of 0.62, indicative of its refined capability to render digital worlds with a higher degree of visual realism and fidelity.
-Mora, with a score of 0.52, although demonstrating a competent level of performance, falls behind Sora, suggesting areas for improvement in achieving the same level of image clarity and detail.
-However, both models achieve identical scores in Appearance Style, at 0.23, which reflects a shared ability to adhere to the stylistic parameters of the digital worlds being simulated.
-This suggests that while there is a difference in the imaging quality, the stylistic translation of textual descriptions into visual aesthetics is accomplished with equivalent proficiency by both models.
 Upon qualitative evaluation, Figure 13 presents samples from Simulate digital worlds tasks, wherein both Sora and Mora were instructed to generated video of "Minecraft" scenes.
@@
 This suggests that Mora is still in a developmental phase, with its generative capabilities requiring further refinement to reach the performance level of Sora.
-5 Discussion
-5.1 Strengths of Mora
-Innovative Framework and Flexibility.
-Mora introduces a groundbreaking multi-agent framework for video generation, significantly advancing the field by enabling a vast array of tasks.
-This innovative approach not only facilitates text-to-video conversion but also supports the simulation of digital worlds, showcasing unparalleled flexibility and efficiency.
-Unlike its closed-source counterparts such as Sora, Mora’s open framework design stands out, offering the ability to seamlessly integrate various models.
-This adaptability allows for the completion of an expanded range of tasks and projects, making Mora an indispensable tool for diverse application scenarios.
-Open-Source Contribution.
-Mora’s open-source nature is highlighted as a significant contribution to the AI community, encouraging further development and refinement by providing a solid foundation upon which future research can build.
-This open-source approach not only democratizes access to advanced video generation technologies but also fosters collaboration and innovation in the field.
-The discussion suggests new ways for future research, including improving the framework’s efficiency, reducing its computational demands during training, and exploring new agent configurations to enhance performance.
-5.2 Limitations of Mora
-Video Dataset is all you need.
-Collecting high-quality video datasets poses significant challenges, primarily due to copyright restrictions on many videos.
-Unlike images, which can often be easier to gather and utilize for training purposes due to a broader range of available resources and more lenient copyright laws, videos frequently contain copyrighted materials that are not as straightforward to collect or use legally.
-This issue is compounded by the fact that high-quality videos often come from professional sources, such as movies, television shows, and proprietary game footage, which are rigorously protected by copyright laws.
-Especially in scenarios involving humans in realistic settings, Mora struggles to generate lifelike movements, such as walking or riding a bicycle.
-Capturing the subtleties of human motion requires not just any video footage, but high-resolution, smoothly captured sequences that detail every aspect of human kinetics, including the nuances of balance, posture, and interaction with surroundings.
-Without access to extensive datasets that accurately represent this wide range of human movements, it becomes challenging for video generation models like Mora to replicate these actions convincingly.
-This limitation underscores the importance of not only the quantity but also the quality and diversity of video datasets in training models to understand and recreate complex human behaviors accurately.
-Quality and Length Gaps.
-Despite its innovative approach, Mora faces notable challenges, especially when compared to its counterpart, Sora, in terms of video generation quality and capabilities.
-While Mora is capable of accomplishing tasks similar to those of Sora, the quality of videos generated by Mora falls significantly short, particularly in scenarios involving substantial object movement.
-Such conditions often introduce a considerable amount of noise, with the quality degradation becoming more pronounced relative to the video length.
-Moreover, although Mora generally maintains quality for up to 12 seconds of video, any attempt to extend beyond this duration leads to a marked decline in video quality.
-In contrast, Sora has demonstrated the ability to produce high-quality videos exceeding one minute in length, as showcased in its technical report.
-This disparity in imaging quality, alongside the limitations in generating videos longer than 12 seconds at a resolution of 1024x576, underscores the urgent need for advancements in Mora’s rendering capabilities and an extension of its video generation parameters.
-Instruction Following Capability.
-Mora has reached a milestone in its development, now capable of generating videos exceeding 10 seconds from given prompts.
-However, despite its ability to include all objects specified in the prompts within the generated videos, Mora encounters limitations in executing certain functions.
-Notably, it struggles with interpreting and rendering the dynamics of motion described in prompts, such as the speed of movement.
-Additionally, Mora lacks the capability to control the direction of motion—specific directions like left or right for specified objects remain unachievable.
-This shortfall primarily stems from the system’s foundational approach to video generation, which operates on an Image to Video basis without direct input from textual prompts.
-Consequently, meeting these specific requirements poses significant challenges, highlighting areas for potential enhancement in Mora’s functionality and interpretative capabilities.
-Human Visual Preference Alignment.
-The absence of human labeling information within the video domain suggests that experimental results may not always align with human visual preferences, highlighting a significant gap.
-For example, during connection video task, the generation of transition videos that disintegrate a male figure only to form a female figure represents an exceedingly illogical scenario.
-This example underscores the necessity for datasets that adhere more closely to physical laws, thereby refining our work and ensuring more realistic and coherent video generation outcomes.
-Such a framework would enhance the appeal of generated videos to viewers and contribute to establishing comprehensive benchmarks and evaluation criteria for the field.
-6 Conclusion
-We introduce Mora, a groundbreaking generalist framework for video generation that addresses a range of video-related tasks.
-Leveraging the collaborative power of multiple agents, Mora marks a considerable advancement in generating videos from textual prompts, establishing new benchmarks for adaptability, efficiency, and output quality in the field of video generation.
-Our thorough evaluation reveals that Mora not only competes with but also exceeds the capabilities of current leading models in certain areas.
-However, it has a notable gap with Sora model by OpenAI whose closed-source nature poses considerable challenges for replication and innovation within the academic and professional communities.
-Our work demonstrates the untapped potential of a meta-programming approach that facilitates intricate collaborations among a variety of agents, each specializing in a segment of the video generation process.
-In essence, the achievements of Mora not only illustrate the current open source state-of-the-art models in video generation but also illuminate the path forward for the field.
-As we continue to explore the vast landscape of generative AI, multi-agent collaboration frameworks like Mora will undoubtedly play a pivotal role in unlocking new creative possibilities and applications, from storytelling and content creation to simulation and training.
-The journey of innovation is far from over, and Mora represents a significant milestone on this ongoing voyage of discovery.
-Looking ahead, there are several promising pathways for further research.
-One such direction involves exploring the integration of more sophisticated natural language understanding capabilities within the agents, potentially allowing for more detailed and context-aware video generations.
-In addition, the expansion of Mora to incorporate real-time feedback loops could offer interactive video creation experiences, where user inputs could guide the generation process in more dynamic and responsive ways.
-Furthermore, the challenge of accessibility and computational resource requirements remains a critical barrier to wider adoption and innovation.
-Future iterations of Mora could benefit from optimizations that reduce these requirements, making advanced video generation technologies more accessible to a broader range of users and developers.
-In parallel, efforts to create more open and collaborative research environments could accelerate progress in this domain, enabling the community to build upon the foundation laid by Mora framework and other pioneering works.
+Figure 14: Samples for text-conditional image-to-video generation of Mora and Sora.
+Prompt for the first line image is: Monster Illustration in flat design style of a diverse family of monsters.
+The group includes a furry brown monster, a sleek black monster with antennas, a spotted green monster, and a tiny polka-dotted monster, all interacting in a playful environment.
+The second image’s prompt is: An image of a realistic cloud that spells "SORA".
+[…]
+Model Video Quality Object Consistency Background Consistency Motion Smoothness Aesthetic Quality Dynamic Degree Imaging Quality Temporal Style
+Mora (Open-Sora-Plan) 0.800 0.98 0.97 0.99 0.66 0.50 0.70 0.31
+Mora (Open-Sora-Plan) 0.767 0.94 0.95 0.99 0.61 0.43 0.68 0.26
+Mora w/o Human-in-the-loop 0.785 0.98 0.96 0.95 0.63 0.50 0.69 0.31
+Mora w/o Self-modulated 0.776 0.96 0.95 0.95 0.62 0.51 0.67 0.27
+Mora with RI-modulated 0.797 0.98 0.96 0.99 0.66 0.50 0.69 0.29
+B.2 Ablation Study
+The results in Table 5 clearly demonstrate the effectiveness of each component in our Mora model.
+The full version of Mora (Open-Sora-Plan) outperforms all ablated variants across most metrics, showcasing the importance of each design choice.
+First, Mora achieves the highest video quality score of 0.800, along with the best object consistency (0.98), background consistency (0.97), motion smoothness (0.99), and imaging quality (0.70).
+This underscores the significance of incorporating all components, as removing any one of them leads to a noticeable drop in performance.
+For instance, removing the human-in-the-loop module (Mora w/o Human-in-the-loop) reduces video quality to 0.785 and slightly decreases motion smoothness (0.95), indicating that human guidance is key to maintaining higher video generation quality and consistency.
+Similarly, removing self-modulation (Mora w/o Self-modulated) results in a further decline in video quality to 0.776, while slightly improving the dynamic degree (0.51).
+However, this increase in dynamic degree comes at the cost of consistency across other metrics, showing that self-modulation balances aesthetic quality and consistency.
+The Random Initial modulation (Mora with RI-modulated) also demonstrates a strong performance, achieving 0.797 in video quality and maintaining the same level of object consistency (0.98) and motion smoothness (0.99) as the full model.
+However, the overall consistency and quality remain lower than the full Mora model, confirming that learned modulation embeddings contribute significantly to the model’s performance.
+These results demonstrate that each component—human-in-the-loop, self-modulation, and optimized modulation embeddings—plays a crucial role in improving the model’s text-to-video generation capabilities.
+The full Mora model’s superior performance validates the importance of our design, providing a clear advantage over the ablated versions.
+Table 6: Framework comparison capabilities of Sora, Mora and other autonomous agent framework. "✓“ indicates the specific feature in the corresponding framework or model. "-“ means absence.
+Framework Role-based agent SOPs Human in the loop Task
+AutoGPT (Aut, ) - - - Code Generation
+AutoGen (Wu et al., 2023b) - Code Generation
+MetaGPT (Hong et al., 2023) - Code Generation
+Sora - - - Video Generation
+Mora Video Generation
+Appendix C Capabilities Analysis
+Compared to the closed-source baseline model Sora and autonomous agents such as MetaGPT (Hong et al., 2023), our framework, Mora, offers enhanced functionalities for video generation tasks.
+As shown in Table 6, Mora encompasses a comprehensive range of capabilities designed to handle diverse and specialized video generation tasks effectively.
+The integration of Standard Operating Procedures (SOPs) such as role-play expertise, structured communication, and streamlined workflows, along with human-in-the-loop systems, significantly refines the control and quality of video generation.
+While other baseline methods can adapt SOP-like designs to boost their performance, they typically do so only within the realm of code generation tasks.
+In contrast, models like Sora lack the capability for fine-grained control or autonomous video generation, highlighting the advanced capabilities of Mora in this domain.
+A detailed comparison of tasks between Mora and Sora is presented in Table 7.
+This comparison demonstrates that, through the collaboration of multiple agents, Mora is capable of accomplishing the video-related tasks that Sora can undertake.
+This comparison highlights Mora’s adaptability and proficiency in addressing a multitude of video generation challenges.
+Table 7: Task comparison between Sora, Mora and other existing models.
+Tasks Example Sora Mora Others
+Text-to-video Generation [Uncaptioned image] (Girdhar et al., 2023; Wang et al., 2023a; Chen et al., 2024a; Ma et al., 2024b)
+Image-to-Video Generation [Uncaptioned image] (Blattmann et al., 2023; pik, ; Gen, a)
+Extend Generated Videos [Uncaptioned image] -
+Video-to-Video Editing [Uncaptioned image] (Molad et al., 2023; Liew et al., 2023; Ceylan et al., 2023)
+Connect Videos [Uncaptioned image] (Chen et al., 2023c)
+Simulate Digital Worlds [Uncaptioned image] -
+Appendix D Discussion
+Advantages of Mora.
+Mora introduces a groundbreaking multi-agent framework for video generation, advancing the field by enabling a variety of tasks such as text-to-video conversion and digital world simulation.
+Unlike closed-source counterparts like Sora, Mora’s open framework offers seamless integration of various models, enhancing flexibility and efficiency for diverse applications.
+As an open-source project, Mora significantly contributes to the AI community by democratizing access to advanced video generation technologies and fostering collaboration and innovation.
+Future research is encouraged to improve the framework’s efficiency, reduce computational demands, and explore new agent configurations to enhance performance.
+Limitations of Mora.
+Mora faces significant limitations, including challenges in collecting high-quality video datasets due to copyright restrictions, resulting in difficulties in generating lifelike human movements.
+Its video quality and length capabilities fall short compared to Sora, with noticeable degradation beyond 12 seconds.
+Mora also struggles with interpreting and rendering motion dynamics from prompts, lacking control over specific directions.
+Furthermore, the absence of human labeling information in video datasets leads to results that may not align with human visual preferences, highlighting the need for datasets that adhere to physical laws.
+Figure 15: Some video examples generated by Mora.